### 07 May 2026 v2

# Some geometric series for Euler’s constant

Abstract.
We provide representations of Euler’s constant $\gamma=0.577$... as series which converge geometrically fast (but use coefficients whose computation induces a quadratic cost). The asymptotic oscillations of these coefficients are discussed.
Key words and phrases:
Euler’s constant, Euler-Mascheroni constant, asymptotic of sequences
2020 Mathematics Subject Classification:
11Y60, 11B83 (Primary) 11B68, 11M41, 40G99 (Secondary)
1. Main result
Let $e_{0}=0$, and define positive rational numbers $e_{m}$ for $m\geq 1$ by this recurrence:
$e_{m}=\frac{2^{m+1}+\sum_{j=1}^{m}\binom{m+1}{j}e_{m-j}}{2^{m+1}-2}.$ (1)
This gives the sequence $2$, $\frac{7}{3}$, $\frac{8}{3}$, $\frac{133}{45}$, $\frac{16}{5}$, … , $e_{10}=\frac{163287}{40579}$, …. See Table 1 for more values.
Table 1. The coefficients $e_{m}$, $1\leq m\leq 20$
$m$ $e_{m}$
11 22
22 $\frac{7}{3}$
33 $\frac{8}{3}$
44 $\frac{133}{45}$
55 $\frac{16}{5}$
66 $\frac{3221}{945}$
77 $\frac{3392}{945}$
88 $\frac{100391}{26775}$
99 $\frac{20848}{5355}$
$10$ $\frac{163287}{40579}$
$m$ $e_{m}$
$11$ $\frac{7567072}{1826055}$
$12$ $\frac{10605587147}{2492565075}$
$13$ $\frac{1551804656}{356080725}$
$14$ $\frac{1732332761353}{388911367845}$
$15$ $\frac{252492267136}{55558766835}$
$16$ $\frac{2313623814645529}{499751107680825}$
$17$ $\frac{261522788700176}{55527900853425}$
$18$ $\frac{69661896931499841923}{14556250513419389775}$
1919 $\frac{2828470111061381408}{582250020536775591}$
2020 $\frac{23101294621895391907711}{4689192129680103420375}$
As usual, $\gamma$ is Euler constant. Here is the main Theorem.
Theorem 1.
Let $\ell\geq 2$. Let $(e_{m})$ be the sequence defined by (1). There holds
$\gamma=\sum_{n=1}^{2^{\ell-1}-1}\frac{1}{n}-(\ell-1)\log 2+\sum_{m=1}^{\infty}\frac{(-1)^{m-1}e_{m}}{m+1}\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{m+1}}.$ (2)
In particular for $\ell=2$:
$\gamma=1-\log 2+\sum_{m=1}^{\infty}\frac{(-1)^{m-1}e_{m}}{m+1}\Bigl(2^{-m-1}+3^{-m-1}\Bigr),$
and for $\ell=3$:
$\gamma=1+\frac{1}{2}+\frac{1}{3}-2\log 2+\sum_{m=1}^{\infty}\frac{(-1)^{m-1}e_{m}}{m+1}\Bigl(4^{-m-1}+5^{-m-1}+6^{-m-1}+7^{-m-1}\Bigr).$
The level $\ell$ is the number of binary digits used by the integers whose inverse powers are added in Equation (2). In Table 2, we illustrate the numerical values obtained using up to $e_{10}$ for $\ell$ being equal to $2$, 33, …, $7$ (so that the computations could have been done by a diligent human calculator), and also up to $e_{20}$ for $\ell=2,3,4$. The computed approximations to Euler’s constant $\gamma$ in this table are truncated not rounded.
Table 2. Table of approximations of $\gamma$
$\ell$ Series from Thm. 1 up to $e_{m}=e_{10}$ last term
2 0.577 15 $\approx\numprint{-0.0001807}$
[…]
$\gamma$ 0.577 215 664 901 532 860 606 512 090
$\ell$ Series from Thm. 1 up to $e_{m}=e_{20}$ last term
[…]
A numerical implementation using Python and the mpmath library is provided at https://burnolmath.gitlab.io/dyadic-gamma/ (it uses by default $\ell=8$). One can express (see Equation (13)) $e_{m}$ as an explicit finite linear combination of the Bernoulli numbers $B_{k}$, $1\leq k\leq m$, but this uses terms which are much larger than $e_{m}$ itself. So, it seems that using the recurrence (1), which is stable numerically, is more appropriate for computing $e_{m}$’s with hundreds of decimal places. However, an algorithm using (1) appears to have an intrinsic (at least) quadratic cost, and is not suited for large scale computations.
[…]
Table 3. Comparison of $e_{m}$ with $(\log 2)^{-1}H_{m+1}$
$m$ $e_{m}-(\log 2)^{-1}H_{m+1}$
11 -0.164 042 561 333
22 -0.311 607 574 963
33 -0.338 948 001 852
44 -0.338 598 121 140
55 -0.334 602 850 177
66 -0.332 236 533 267
77 -0.331 621 032 426
88 -0.331 908 031 990
99 -0.332 424 034 678
$10$ -0.332 833 436 510
$m$ $e_{m}-(\log 2)^{-1}H_{m+1}$
$11$ -0.333 041 375 341
$12$ -0.333 074 427 538
$13$ -0.332 999 205 051
$14$ -0.332 880 028 058
$15$ -0.332 762 836 214
$16$ -0.332 673 218 057
$17$ -0.332 620 451 671
$18$ -0.332 603 053 818
1919 -0.332 613 748 636
2020 -0.332 643 144 549
Figure 1. $e_{m}-(\log 2)^{-1}H_{m+1}$ vs $\log_{2}(m)$, $16\leq m\leq 500$
Theorem 2.
There holds
$e_{m}=\log_{2}(m)+\frac{1}{2}+\phi(\log_{2}(m))+O(\frac{\log m}{m})$ (3)
with
$\phi(t)=-(\log 2)^{-1}\sum_{n\in\mathbb{Z},n\neq 0}\Gamma\bigl(1-\frac{2\pi in}{\log 2}\bigr)\zeta\bigl(1-\frac{2\pi in}{\log 2}\bigr)\mathsf{e}^{2\pi int}.$
[…]
Figure 2. $e_{m}-(\log 2)^{-1}H_{m+1}$ as function of $\log_{2}(m)$, and the graph of $\frac{\gamma}{\log 2}+\frac{1}{2}+\phi(t)$, $16\leq m\leq 1000$
As a corollary to the Theorem, $e_{2m}\approx 1+e_{m}$, a fact which does not seem immediately obvious from Equation 1.
Remark 1.
Numerical investigations suggest that the logarithm in the numerator of the error term from Equation (3) can be omitted. However we should remain cautious because we can compute in reasonable time only of the order of ten thousand coefficients. To be more precise, the error term appears to have the shape $\psi(\log_{2}(m))/m+o(1/m)$ with some $1$-periodic function $\psi$ averaging to $\frac{3}{2\log(2)}$.
Remark 2.
[…]
Remark 3.
We will establish concrete explicit upper and lower bounds for $e_{m}$. They are useful, especially the upper bound, to calibrate numerical implementations.
2. Geometric series for $\gamma$
We obtain Theorem 1 as a corollary to our previous study of the Euler alternating series [5].
Proof.
In [5, Eq. (5)], a sequence of meromorphic functions $(c_{m}(s))$ , $m\geq 0$, $s\in\mathbb{C}$, is defined by the initial condition $c_{0}(s)=1$ and, for $m\geq 1$:
$c_{m}(s)=\frac{1}{2^{m+s}-2}\sum_{j=1}^{m}\binom{m}{j}c_{m-j}(s).$ (4)
In [5], there is an integer parameter $b>1$, which here has been set to $b=2$. There are generalizations of Theorem 1 to general $b>2$, where the analog of the recurrence from Equation (1) has additional factors given by the power sums of the base-$b$ digits.
Let $\ell\geq 2$, the level, be some integer greater than one. The Euler alternating series $\eta(s)=\sum_{n=1}^{\infty}(-1)^{n-1}n^{-s}$ was proven in [5, Thm. 1] to verify for $\Re s>0$ the next formula, which converges locally uniformly:
$\eta(s)=\frac{2^{s}-2}{2^{s}}\sum_{0<n<2^{\ell-1}}\frac{1}{n^{s}}+\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{s}}+\sum_{m=1}^{\infty}(-1)^{m}\frac{(s)_{m}}{m!}c_{m}(s)\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{s+m}}\,.$ (5)
We now compute the derivative at $s=1$ (here $H_{0}=0$):
$\eta^{\prime}(1)=(\log 2)H_{2^{\ell-1}-1}+\sum_{m=0}^{\infty}(-1)^{m}\left\{\begin{aligned} &H_{m}c_{m}(1)\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{m+1}}\\ &+c_{m}^{\prime}(1)\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{m+1}}\\ &-c_{m}(1)\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{\log n}{n^{m+1}}\end{aligned}\right$. (6)
Recall from [5] that $c_{m}(1)=(m+1)^{-1}$ (which is easily checked on the recurrence (4)). Leaving temporarily aside the consideration of the $c_{m}^{\prime}(1)$, we compute:
$\sum_{m=0}^{\infty}(-1)^{m}\biggl(\frac{-1}{m+1}\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{\log n}{n^{m+1}}+\frac{H_{m}}{m+1}\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{m+1}}\biggr)$
$=\sum_{2^{\ell-1}\leq n<2^{\ell}}\left(-\log n\log(1+\frac{1}{n})+\left.\frac{\mathrm{d}}{\mathrm{d}s}\right|_{s=1}\sum_{m=1}^{\infty}\frac{(-1)^{m}}{m+1}\frac{(s)_{m}}{m!}(1/n)^{m+1}\right)$
$=\sum_{2^{\ell-1}\leq n<2^{\ell}}\left(-\log n\log(1+\frac{1}{n})+\left.\frac{\mathrm{d}}{\mathrm{d}s}\right|_{s=1}\int_{0}^{\frac{1}{n}}(1+x)^{-s}\mathrm{d}x\right)$
$=\sum_{2^{\ell-1}\leq n<2^{\ell}}\left(-\log n\log(1+\frac{1}{n})+\int_{0}^{\frac{1}{n}}\frac{-\log(1+x)}{(1+x)}\mathrm{d}x\right)$
$=\sum_{2^{\ell-1}\leq n<2^{\ell}}\left(-\log n\log(1+\frac{1}{n})-\frac{1}{2}\log^{2}(1+\frac{1}{n})\right)$
$=-\frac{1}{2}\sum_{2^{\ell-1}\leq n<2^{\ell}}\log(1+\frac{1}{n})\log\bigl(n^{2}(1+\frac{1}{n})\bigr)=-\frac{1}{2}\sum_{2^{\ell-1}\leq n<2^{\ell}}\bigl(\log^{2}(n+1)-\log^{2}(n)\bigr)$
$=-\frac{1}{2}\Bigl(\ell^{2}\log^{2}2-(\ell-1)^{2}\log^{2}2\Bigr)=-(\ell-\frac{1}{2})\log^{2}2.$
Hence, from Equation (6):
$\eta^{\prime}(1)=(\log 2)H_{2^{\ell-1}-1}-(\ell-\frac{1}{2})\log^{2}2+\sum_{m=0}^{\infty}(-1)^{m}c_{m}^{\prime}(1)\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{m+1}}.$
Now, as is well-known, $\eta^{\prime}(1)=\log(2)(\gamma-\frac{1}{2}\log 2)$, so, with the definition
$e_{m}=-(m+1)\frac{c_{m}^{\prime}(1)}{\log 2},$ (7)
(in particular $e_{0}=0$), we obtain
$\gamma=H_{2^{\ell-1}-1}-(\ell-1)\log 2+\sum_{m=1}^{\infty}(-1)^{m-1}\frac{e_{m}}{m+1}\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{m+1}}.$
The proof of Theorem 1 will be complete once it has been confirmed that the $(e_{m})$ sequence defined by Equation (7) verifies the recurrence (1). Taking the derivative at $s=1$ of Equation (4), and recalling $c_{m}(1)=(m+1)^{-1}$, we obtain, for $m\geq 1$:
$c_{m}^{\prime}(1)$ $=\frac{-(\log 2)2^{m+1}}{2^{m+1}-2}c_{m}(1)+(2^{m+1}-2)^{-1}\sum_{j=1}^{m}\binom{m}{j}c_{m-j}^{\prime}(1)$
$e_{m}$ $=\frac{2^{m+1}}{2^{m+1}-2}+(2^{m+1}-2)^{-1}\sum_{j=1}^{m}\binom{m}{j}\frac{m+1}{m-j+1}e_{m-j}.$
This, indeed, is Equation (1). And $e_{0}=0$. ∎
3. Estimates for the coefficients $e_{m}$
In this section, we establish the next proposition as a corollary to the recurrence from Equation (1).
Proposition 1.
For $m\geq 2$, there holds
$\frac{H_{m+1}}{\log 2}-\numprint{0.35}<e_{m}<\frac{H_{m+1}}{\log 2}-\numprint{0.31}\;.$
We first establish a less tight bound. We will then re-use the tools from the proof of the Lemma 1 to prove Proposition 1.
Lemma 1.
There holds for any $m\in\mathbb{N}$ (with $H_{n}=\sum_{j=1}^{n}j^{-1}$):
$\frac{H_{m+1}-1}{\log 2}\leq e_{m}<\frac{H_{m+1}}{\log 2}-\numprint{0.161}\,.$
Proof.
Let us define $\delta_{n}=e_{n}-(\log 2)^{-1}H_{n+1}$ for $n\geq 0$. We thus have, for $m\geq 1$:
$(2^{m+1}-2)e_{m}=\left\{\begin{aligned} 2^{m+1}&+(\log 2)^{-1}\sum_{j=1}^{m}\binom{m+1}{j}H_{m-j+1}\\ &+\sum_{j=1}^{m}\binom{m+1}{j}\delta_{m-j}.\end{aligned}\right$. (8)
Now (as is well-known):
$H_{m+1}+\sum_{j=1}^{m}\binom{m+1}{j}H_{m-j+1}$ $=\sum_{k=1}^{m+1}\binom{m+1}{k}H_{k}$
$=\int_{0}^{1}\sum_{k=0}^{m+1}\binom{m+1}{k}\frac{1-t^{k}}{1-t}\mathrm{d}t$
$=\int_{0}^{1}\frac{2^{m+1}-(1+t)^{m+1}}{1-t}\mathrm{d}t$
$=\int_{1}^{2}\frac{2^{m+1}-u^{m+1}}{2-u}\mathrm{d}u$
$=2^{m+1}\int_{\frac{1}{2}}^{1}\frac{1-x^{m+1}}{1-x}\mathrm{d}x$
$=2^{m+1}H_{m+1}-2^{m+1}\int_{0}^{\frac{1}{2}}\frac{1-x^{m+1}}{1-x}\mathrm{d}x$
$=2^{m+1}H_{m+1}-2^{m+1}\log 2+2^{m+1}\underbrace{\int_{0}^{\frac{1}{2}}\frac{x^{m+1}}{1-x}\mathrm{d}x}_{\epsilon_{m}}.$
Hence, from (8),
$(2^{m+1}-2)e_{m}=(2^{m+1}-2)\frac{H_{m+1}}{\log 2}+\frac{H_{m+1}}{\log 2}+2^{m+1}\frac{\epsilon_{m}}{\log 2}+\sum_{j=1}^{m}\binom{m+1}{j}\delta_{m-j}.$
$e_{m}=\frac{H_{m+1}}{\log 2}+\underbrace{\frac{H_{m+1}+2^{m+1}\epsilon_{m}}{(2^{m+1}-2)\log 2}+\frac{\sum_{j=1}^{m}\binom{m+1}{j}\delta_{m-j}}{2^{m+1}-2}}_{\delta_{m}}$ (9)
Suppose now by induction that $\delta_{j}\geq-(\log 2)^{-1}$, for $0\leq j<m$, which is true for $m=1$, as $\delta_{0}=-(\log 2)^{-1}$. The above implies $\delta_{m}\geq-(\log 2)^{-1}$ too, so it is true for all $m$.
To obtain an upper bound, we define
$\mu_{m}=\max_{0\leq j\leq m}\delta_{j}.$
In particular $\mu_{0}=\delta_{0}=-(\log 2)^{-1}\approx\numprint{-1.443}$, and $\mu_{1}=\delta_{1}=2-(\log 2)^{-1}\frac{3}{2}\approx\numprint{-0.164}$. The above implies, for $m\geq 1$:
$\mu_{m}\leq\frac{H_{m+1}+2^{m+1}\epsilon_{m}}{(2^{m+1}-2)\log 2}+\mu_{m-1},$
hence by induction, we obtain, for any given $m$:
$p>m$ $\implies\mu_{p}\leq\mu_{m}+(\log 2)^{-1}\sum_{n=m+1}^{p}\frac{H_{n+1}+2^{n+1}\epsilon_{n}}{2^{n+1}-2},$
thus, using
$\frac{1}{2^{n+1}-2}$ $=\frac{1}{2^{n+1}}+\frac{2}{2^{n+1}(2^{n+1}-2)}\leq\frac{1}{2^{n+1}}+\frac{1}{4^{n}},$
we obtain, for any $p>m\geq 0$:
$\delta_{p}\leq\mu_{m}+(\log 2)^{-1}\sum_{n=m+1}^{\infty}\left(2^{-n-1}H_{n+1}+4^{-n}H_{n+1}+\epsilon_{n}+2^{-n+1}\epsilon_{n}\right).$
Now,
$\sum_{n=m+1}^{\infty}2^{-n-1}H_{n+1}=\frac{-\log(1-\frac{1}{2})}{1-\frac{1}{2}}-\sum_{n=0}^{m}2^{-n-1}H_{n+1},$
and
$\sum_{n=m+1}^{\infty}4^{-n}H_{n+1}=4\frac{-\log(1-\frac{1}{4})}{1-\frac{1}{4}}-\sum_{n=0}^{m}4^{-n}H_{n+1}.$
Evaluating numerically for $m=10$ we obtain respectively:
$(\log 2)^{-1}\sum_{n=11}^{\infty}2^{-n-1}H_{n+1}$ $\approx\numprint{0.002237}$
[…]
We bound trivially $\epsilon_{n}=\int_{0}^{\frac{1}{2}}x^{n+1}(1-x)^{-1}\mathrm{d}x<2^{-n-1}$, hence:
$(\log 2)^{-1}\sum_{n=11}^{\infty}(1+2^{-n+1})\epsilon_{n}$ $<(\log 2)^{-1}(1+2^{-10})\sum_{n=11}^{\infty}2^{-n-1}$
$=(\log 2)^{-1}(2^{-11}+2^{-21})\approx\numprint{0.000705}$
So for $p>10$, $\delta_{p}<\mu_{10}+\numprint{0.003}$. A quick look at Table 3 shows that $\mu_{10}=\delta_{1}=2-(\log 2)^{-1}\frac{3}{2}\approx\numprint{-0.164043}$. Hence $\mu_{10}+\numprint{0.003}<\numprint{-0.161}$. This concludes the proof of Lemma 1. ∎
Proof of Proposition 1.
We continue with the same technique as in the previous proof. To improve the upper bound, we separate out $\delta_{1}$:
$(2^{m+1}-2)^{-1}\sum_{j=1}^{m}\binom{m+1}{j}\delta_{m-j}$ $\leq\frac{m(m+1)}{2(2^{m+1}-2)}\delta_{1}$
$\mkern-100.0mu+\frac{2^{m+1}-2-m(m+1)/2}{2^{m+1}-2}\max\{\delta_{j},0\leq j<m,j\neq 1\}.$
As we know that $\delta_{j}\geq\delta_{0}=-(\log 2)^{-1}$, we can further bound this as
$\leq\frac{m(m+1)}{2(2^{m+1}-2)}\bigl(\underbrace{\delta_{1}+|\delta_{0}|}_{2-\frac{1}{2}(\log 2)^{-1}\approx\numprint{1.279}}\bigr)+\max\{\delta_{j},0\leq j<m,j\neq 1\}.$
We again use $(2^{n+1}-2)^{-1}<2^{-n-1}(1+2^{-n+1})$ for $n\geq 1$. So:
$\sum_{n=21}^{\infty}\frac{n(n+1)}{2(2^{n+1}-2)}$ $<(1+2^{-20})\frac{1}{8}\sum_{n=21}^{\infty}\frac{n(n+1)}{2^{n-1}}$
$=(1+2^{-20})\frac{1}{8}\left(\frac{2}{(1-\frac{1}{2})^{3}}-\sum_{k=0}^{20}\frac{k(k+1)}{2^{k-1}}\right)$
$\approx\numprint{0.000121117}$
The product of this with $\delta_{1}-\delta_{0}$ is less than $\numprint{0.0002}$. We still need to take into account the contribution to the upper bound from the series with the $H_{n+1}$’s and the $\epsilon_{n}$’s. We expect something about one thousand times smaller than what we computed in the previous proof for the remainders, and indeed, one obtains:
[…]
$(\log 2)^{-1}\sum_{n=21}^{\infty}(1+2^{-n+1})\epsilon_{n}$ $<(\log 2)^{-1}(1+2^{-20})\sum_{n=21}^{\infty}2^{-n-1}$
[…]
As $\delta_{2}\approx\numprint{-0.311608}$ is, as we see from Table 3, the maximum among the $\delta_{m}$’s, for $m\neq 1$ and $0\leq m\leq 20$, we conclude, with some margin, that
$m\geq 2\implies\delta_{m}<\numprint{-0.3114}\;.$
Similarly, regarding the lower bound, we can separate out $\delta_{0}=-(\log 2)^{-1}$, writing
$(2^{p+1}-2)^{-1}\sum_{j=1}^{p}\binom{p+1}{j}\delta_{p-j}\geq\frac{p+1}{(2^{p+1}-2)}\delta_{0}+\min\{\delta_{j},1\leq j<p\},$
where we used the fact that we know $\delta_{j}\leq 0$ for all $j$’s. Now working again with $p>10$, we obtain (as the other contributions involving the $H_{n+1}$’s and the $\epsilon_{n}$’s are positive):
$p>10\implies\delta_{p}\geq\delta_{0}\sum_{n=11}^{\infty}\frac{n+1}{(2^{n+1}-2)}+\underbrace{\min\{\delta_{j},1\leq j\leq 10\}}_{{}=\delta_{3}}.$
Now
$\sum_{n=11}^{\infty}\frac{n+1}{(2^{n+1}-2)}$ $<(1+2^{-10})\sum_{n=11}^{\infty}\frac{n+1}{2^{n+1}}$
$=(1+2^{-10})\frac{1}{2}\left(\frac{1}{(1-\frac{1}{2})^{2}}-\sum_{k=0}^{10}(k+1)2^{-k}\right)$
$\approx\numprint{0.006354}.$
The product of this with $\delta_{0}$ is about $\numprint{-0.009167}>\numprint{-0.01}$ and we thus can conclude that for all $m\geq 1$, $\delta_{m}>\numprint{-0.01}+\delta_{3}>\numprint{-0.35}$. ∎
4. The asymptotic behavior of $(e_{m})$
We have seen in the proof of Theorem 1 that we can define $e_{m}$ as $(m+1)c_{m}^{\prime}(1)/\log 2$ where the meromorphic functions $c_{m}(s)$ verify the recurrence Equation (4). From [5] and [4], there holds $c_{m}(s)=\frac{2^{s}-2}{2^{s}}u_{m}(s)$, where the meromorphic function $u_{m}(s)$ is, for $\Re s>1$, the $m$th moment of a certain discrete (complex) measure on $[0,1]$:
$\Re s>1\implies u_{m}(s)=\sum_{l=0}^{\infty}\biggl(\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)2^{-ls}\,.$ (10)
In particular $u_{0}(s)=\sum_{l=0}^{\infty}2^{-l(s-1)}=2^{s}/(2^{s}-2)$. As the function $t\mapsto t^{m}$ is non-decreasing and convex, there holds
$0\leq\int_{0}^{1}t^{m}\mathrm{d}t-2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\leq\frac{1}{2}2^{-l}.$ (11)
So, at first for $\Re s=1$:
$u_{m}(s)$ $=(m+1)^{-1}\sum_{l=0}^{\infty}2^{-l(s-1)}-\sum_{l=0}^{\infty}\biggl((m+1)^{-1}-2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}$
$(m+1)u_{m}(s)$ $=\frac{2^{s}}{2^{s}-2}-\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}$
$(m+1)c_{m}(s)$ $=1-\frac{2^{s}-2}{2^{s}}\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}.$
Thanks to estimate (11), this gives the analytic continution to $\Re s>0$, and we can now compute the value of the derivative at $s=1$:
$(m+1)c_{m}^{\prime}(1)=-\log(2)\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr).$
From Equation (7) which arose in the course of the proof of Theorem 1, we thus get an “explicit” formula for the coefficients $e_{m}$:
Proposition 2.
There holds:
$e_{m}=\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)\,.$ (12)
In terms of Bernoulli polynomials and numbers:
$e_{m}=\sum_{l=0}^{\infty}\biggl(1-\frac{B_{m+1}(2^{l})-B_{m+1}(0)}{2^{(m+1)l}}\biggr)=-\sum_{k=1}^{m}\binom{m+1}{k}B_{k}\frac{2^{k}}{2^{k}-1}\,.$ (13)
Remark 4.
The Bernoulli polynomials and numbers used in (13) are defined by $\sum_{n=0}^{\infty}B_{n}(x)\frac{t^{n}}{n!}=\frac{t\mathsf{e}^{xt}}{\mathsf{e}^{t}-1}$, $B_{n}(x)=\sum_{k=0}^{n}\binom{n}{k}B_{k}x^{n-k}$, $B_{0}=1$, $B_{1}=-\frac{1}{2}$, $B_{n}=B_{n}(0)$, $(m+1)\sum_{0\leq n<N}n^{m}=B_{m+1}(N)-B_{m+1}(0)$. Formula (13) has large contributions of both signs and does not look promising numerically. In the sequel we will only use (12).
Remark 5.
A variant of (13) is (see Equation (22) for the Euler polynomials):
$e_{m}$ $=\sum_{l=0}^{\infty}(l+1)\frac{B_{m+1}(2^{l+1})-2^{m+1}B_{m+1}(2^{l})+(2^{m+1}-1)B_{m+1}(0)}{2^{(l+1)(m+1)}}$
$=\frac{m+1}{2}\sum_{l=0}^{\infty}(l+1)\frac{E_{m}(2^{l+1})-E_{m}(0)}{2^{(l+1)(m+1)}}\,.$
One only needs to write the numerator as $(l+1)(v_{l+1}-2^{m+1}v_{l})$ with $v_{l}=B_{m+1}(2^{l})-B_{m+1}(0)-2^{(m+1)l}$ and rearrange the sum. Details are left to the reader.
We will derive in this section from Equation (12) some asymptotic for $e_{m}$ as $m\to\infty$. From (11), the $l$th term is non-negative and bounded above by $(m+1)2^{-l-1}$. Let $L_{m}=\lfloor 2\log_{2}(m+1)\rfloor$. The sum of the contributions for $l>L_{m}$ is bounded above by $(m+1)2^{-L_{m}-1}$, hence, by $(m+1)^{-1}$.
To approximate the ($l$ dependent) moments $p_{l}=2^{-l}\sum_{0\leq n<2^{l}}(2^{-l}n)^{m}$ for $l\leq L_{m}$, we shall apply the same technique as in [6, Prop. 1]. For the convenience of the reader, the presentation here is entirely self-contained. [^1] In v1 of [6], which is the current version at time of writing, the Proposition 1 is formulated in a way which does not make explicit the implied constants in the big-O, although this is quite easy to extract from the proof there. As here we need to apply this to mutliple probability measures, we had anyhow to reproduce parts of the argument from [6] in order to control the uniformity. Certainly, the published version of [6] will make more explicit the big-O bound. Let us define
$q_{l}=2^{-l}\sum_{0\leq n<2^{l}}\mathsf{e}^{-m(1-2^{-l}n)}=\frac{1-\mathsf{e}^{-m}}{2^{l}(\mathsf{e}^{m2^{-l}}-1)}\,.$ (14)
There holds $0\leq x\mathsf{e}^{-x}\leq\mathsf{e}^{-1}$ for $x\geq 0$, hence $x^{m}\leq\mathsf{e}^{-m(1-x)}$. In particular $(2^{-l}n)^{m}\leq\mathsf{e}^{-m(1-2^{-l}n)}$ and $0\leq p_{l}\leq q_{l}$. We split $q_{l}-p_{l}$ into two parts:
$q_{l}-p_{l}=2^{-l}(\sum_{0\leq n<2^{l-1}}+\sum_{2^{l-1}\leq n<2^{l}})\biggl(\mathsf{e}^{-m(1-2^{-l}n)}-(2^{-l}n)^{m}\biggr)$ (15)
We bound trivially the first part by
$2^{-l}\sum_{0\leq n<2^{l-1}}\mathsf{e}^{-m(1-2^{-l}n)}\leq\int_{0}^{\frac{1}{2}}\mathsf{e}^{m(x-1)}\mathrm{d}x=\frac{1}{m}\mathsf{e}^{-m/2}(1-\mathsf{e}^{-m/2}).$
This, and everything following applies also to non-integer positive $m$’s, the quantity $e_{m}$ can be defined for them using (12).
Regarding the second part in Equation (15), we use (as in the proof of [6, Prop. 1]) the bound
$m\geq 0,\;\frac{1}{2}\leq x\leq 1\implies\mathsf{e}^{-m(1-x)}-x^{m}\leq m(1-x)^{2}\mathsf{e}^{-m(1-x)}.$
So, for positive $m$:
$0\leq q_{l}-p_{l}\leq\frac{\mathsf{e}^{-m/2}}{m}+2^{-l}\sum_{2^{l-1}\leq n<2^{l}}m(1-2^{-l}n)^{2}\mathsf{e}^{-m(1-2^{-l}n)}$
(Note that if $l=0$, the sum above is empty and has value 0; so we now assume $l\geq 1$ and the bound we obtain also applies with $l=0$). Let $\psi_{m}(x)=\frac{\mathrm{d}}{\mathrm{d}x}\left((1-x)^{2}\mathsf{e}^{-m(1-x)}\right)=\bigl(m(1-x)-2\bigr)(1-x)\mathsf{e}^{-m(1-x)}$. We write:
$2^{-l}\sum_{2^{l-1}\leq n<2^{l}}m(1-2^{-l}n)^{2}\mathsf{e}^{-m(1-2^{-l}n)}=\frac{m}{8}\mathsf{e}^{-m/2}+m2^{-l}\sum_{2^{l-1}\leq n<2^{l}}\int_{\frac{1}{2}}^{2^{-l}n}\psi_{m}(y)\mathrm{d}y$
Using the probability measure $\mu_{l}=2^{-l}\sum_{0\leq n<2^{l}}\delta_{2^{-l}n}$, which has the property $\mu_{l}([y,1])\leq 1-y$ for every $y\in[0,1]$, we can rewrite the second summand on the right-hand side as:
$)\displaystyle m\int_{[\frac{1}{2},1\rparen}\int_{\frac{1}{2}}^{x}\psi_{m}(y)\mathrm{d}y\,\mu_{l}(\mathrm{d}x$) $\leq m\int_{\frac{1}{2}\leq y\leq x<1}\max(\psi_{m}(y),0)\mathrm{d}y\,\mu_{l}(\mathrm{d}x)$
$\leq m\int_{\frac{1}{2}}^{1}\max(\psi_{m}(y),0)(1-y)\mathrm{d}y$
$=m\int_{\frac{1}{2}}^{\max(1-\frac{2}{m},\frac{1}{2})}\bigl(m(1-y)-2\bigr)(1-y)^{2}\mathsf{e}^{-m(1-y)}\mathrm{d}y$
$=m^{-2}\int_{\min(2,m/2)}^{m/2}(t-2)t^{2}\mathsf{e}^{-t}\mathrm{d}t\leq 18\mathsf{e}^{-2}m^{-2}.$
We have obtained, for positive $m$ (with some loss of precision near zero, but we are only interested into $m\to\infty$):
$0\leq q_{l}-p_{l}\leq(m^{-1}+\frac{m}{8})\mathsf{e}^{-m/2}+\frac{18}{\mathsf{e}^{2}m^{2}}\leq C_{1}m^{-2},$ (16)
with $C_{1}=\max_{t\geq 0}(t+t^{3}/8)\mathsf{e}^{-t/2}+18\mathsf{e}^{-2}$. It turns out that the maximum is attained at $t\approx\numprint{5.04}$ and its value is $\approx\numprint{1.693}$ and $18\mathsf{e}^{-2}\approx\numprint{2.436}$ and $C_{1}<\numprint{4.13}$. Of course when $m$ is large $18\mathsf{e}^{-2}+\epsilon$ gives a tighter bound.
The proof applies to any positive $m$, and for $m$ a positive integer it becomes a statement involving the values of Bernoulli polynomials at $N=2^{l}$. That $N$ was a power of $2$ is not important. We formulate here a dedicated statement, as we could not find such property in the standard references.
Proposition 3.
For any two positive integers $m$ and $n$, the following uniform estimate holds:
$0\leq\frac{1-\mathsf{e}^{-m}}{n\bigl(\mathsf{e}^{mn^{-1}}-1\bigr)}-\frac{B_{m+1}(n)-B_{m+1}(0)}{(m+1)n^{m+1}}\leq\frac{5}{m^{2}}.$ (17)
Proof.
Let’s use $N$ rather than $n$ which has served earlier as summation index. Recall $\sum_{0\leq n<N}n^{m}=(m+1)^{-1}\bigl(B_{m+1}(N)-B_{m+1}(0)\bigr)$. The proof leading to (16) works identically with an arbitrary positive integer $N$ in the place of $2^{l}$, if $N$ is even. If $N\geq 3$ is odd, $N=2M+1$, $M\geq 1$, we have to be careful about how we split (the analog of) $q_{l}-p_{l}$ into two parts. Formerly we compared the first part with an integral from 0 to $\frac{1}{2}$ but now, using $0\leq n<\frac{1}{2}N$, i.e., $0\leq n\leq M$, we see that the rectangle basis extends a bit further to $\frac{1}{2}+\frac{1}{2N}$. We could improve that (for $m$ large) by bounding the contribution of $x=M/N$ pointwise by $\mathsf{e}^{-m/2}$ (as done in [6, Prop. 1] which handles general positive measures). But let us continue with $\int_{0}^{\frac{1}{2}+\frac{1}{2N}}\mathsf{e}^{m(x-1)}\mathrm{d}x<\mathsf{e}^{m/2N}m^{-1}\mathsf{e}^{-m/2}\leq\mathsf{e}^{m/6}m^{-1}\mathsf{e}^{-m/2}$, for the first part. For the second part, i.e. for the contributions of the points $x=N^{-1}n$, $M+1\leq n<N$, we can still use $x=\frac{1}{2}$ as reference point, there is no change, except that we could gain a little in the $\frac{m}{8}\mathsf{e}^{-m/2}$ term (as the weight of the half-interval is now not $1/2$, but $M/N$). Anyway, we end up with a slightly deteroriated final global upper bound $\mathsf{e}^{m/6}m^{-1}\mathsf{e}^{-m/2}+\frac{m}{8}\mathsf{e}^{-m/2}+18\mathsf{e}^{-2}m^{-2}$, which is $\leq C_{2}m^{-2}$ with $C_{2}=\max_{t\geq 0}(t\mathsf{e}^{-t/3}+\frac{t^{3}}{8}\mathsf{e}^{-t/2})+18\mathsf{e}^{-2}$. The maximum is again attained near $t=5$, it is found to be $\numprint{2.27}$ and $C_{2}<4.67$. ∎
As a corollary if we let both $m$ and $n$ go to infinity such that the ratio converges to some value $q$, we get (with $1$ on the right if $q=0$):
$\lim_{m,n\to\infty,mn^{-1}\to q}\frac{B_{m+1}(n)-B_{m+1}(0)}{n^{m+1}}=\frac{q}{\mathsf{e}^{q}-1}\,.$
This is indeed what one would expect from expressing the Bernoulli polynomials in terms of Bernoulli numbers and taking the limit term-wise, but (17) makes it easy (and gives an $O(m^{-1})$ error estimate). Reference [11], which uses the saddle-point method in the complex domain, provides full asymptotics for, among others, $B_{n}(nz+\frac{1}{2})$ as $n\to\infty$.
We now resume the study of the quantities $e_{m}$. Recall that we have set $L_{m}=\lfloor 2\log_{2}(m+1)\rfloor$ and already estimated
$0\leq\sum_{l>L_{m}}\bigl(1-(m+1)p_{l}\bigr)\leq\frac{1}{m+1},$
so, combining this with Equations (12) and (16):
$\sum_{0\leq l\leq L_{m}}\bigl(1-(m+1)q_{l}\bigr)\leq e_{m}\leq\sum_{0\leq l\leq L_{m}}\bigl(1-(m+1)q_{l}\bigr)+(L_{m}+1)(m+1)\frac{C_{1}}{m^{2}}+\frac{1}{m+1}\,.$
Let us recall Equation (14) defining $q_{l}$ and record our current result:
$e_{m}=\sum_{0\leq l\leq L_{m}}\bigl(1-(m+1)\frac{1-\mathsf{e}^{-m}}{2^{l}(\mathsf{e}^{m2^{-l}}-1)}\bigr)+O(\frac{\log m}{m}).$ (18)
This motivates the definition of auxiliary coefficients $f_{m}$ and $g_{m}$, for positive real $m$:
$f_{m}$ $=\sum_{l=0}^{\infty}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)$ (19)
$g_{m}$ $=-\sum_{l=-1}^{-\infty}\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}$ (20)
Clearly the series defining $f_{m}$ converges uniformly for $m$ bounded. We bound trivially the terms in the series defining $g_{m}$::
$x>0\implies\frac{x}{\mathsf{e}^{x}-1}=\frac{x\mathsf{e}^{-x/2}}{2\sinh(x/2)}\leq\mathsf{e}^{-x/2}$
$0\leq-g_{m}\leq\frac{2m}{\mathsf{e}^{2m}-1}+\sum_{n=4}^{\infty}\mathsf{e}^{-nm/2}=O_{m\to\infty}(me^{-2m}).$
[…]
$\phi(t)=-t-\frac{1}{2}+\sum_{l=0}^{\infty}\bigl(1-\frac{2^{t-l}}{\mathsf{e}^{2^{t-l}}-1}\bigr)-\sum_{l=-1}^{-\infty}\frac{2^{t-l}}{\mathsf{e}^{2^{t-l}}-1}\,.$ (21)
The reason for the extra term $\frac{1}{2}$ is that this makes the average of $\phi$ vanish, as will be seen later. As shown previously the two series are uniformly convergent for $0\leq t\leq 1$, so $\phi$ is a continuous function. Here is a stronger (elementary) statement:
Proposition 4.
Equation (21) defines $\phi(t)$ as an analytic function of $t$ in the horizontal strip $|\Im t|<\frac{\pi}{2\log(2)}$.
Proof.
Details are left to reader: $\sum_{l=0}^{\infty}(1-\frac{2^{-l}z}{\mathsf{e}^{2^{-l}z}-1})$ defines a meromorphic function in the entire complex plane, with poles at the elements of $2\pi i\mathbb{Z}\setminus\{0\}$, so in particular if we substitute $z=2^{t}$, we obtain an analytic function in the strip $|\Im t|<\frac{\pi}{2\log 2}$. Regarding $\sum_{k=1}^{\infty}\frac{2^{k}2^{t}}{\mathsf{e}^{2^{k}2^{t}}-1}$, it is absolutely and uniformly convergent if $a<\Re t<b$, $|\Im t|<\frac{\pi}{2}-\epsilon$, $0<\epsilon<\frac{\pi}{2}$, from bounding below the denominator and above the numerator. This proves that $\phi(t)$ is as stated. ∎
Returning to Equation (18), we get:
$e_{m}$ $=L_{m}+1-\frac{m+1}{m}(1-\mathsf{e}^{-m})\biggl(L_{m}+1-\sum_{0\leq l\leq L_{m}}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)\biggr)+O(\frac{\log m}{m})$
$=(1+O(\frac{1}{m}))\Bigl(f_{m}-\sum_{l>L_{m}}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)\Bigr)+O(\frac{\log m}{m}).$
For $l>L_{m}=\lfloor 2\log_{2}(m+1)\rfloor$, $2^{-l}m\leq m/(m+1)^{2}$, hence:
$\sum_{l>L_{m}}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)=O(\sum_{l>L_{m}}\frac{m^{2}}{4^{l}})=O(\frac{m^{2}}{4^{L_{m}}})=O(m^{-2}).$
This gives, for $m\to+\infty$:
$e_{m}=(1+O(\frac{1}{m}))\Bigl(\log_{2}(m)+\frac{1}{2}+\phi(\log_{2}(m))+O(m^{-2})+O(m\mathsf{e}^{-2m})\Bigr)+O(\frac{\log m}{m}).$
As $\phi$ is periodic, it is bounded and we have obtained the proof of the main estimate as given in Theorem 2:
Proposition 5.
The quantities $(e_{m})$ defined by the recurrence (1), or, more generally, defined for all real positive $m$ by Equation (12), verify the asymptotic estimate, as $m\to+\infty$:
$e_{m}=\log_{2}(m)+\frac{1}{2}+\phi(\log_{2}(m))+O(\frac{\log m}{m}),$
where $\phi(t)$ is the function defined by Equation (21). It is 11-periodic and analytic for $|\Im t|<\frac{\pi}{2\log(2)}$.
We prove in the next section that $\phi(t)$ has zero mean.
5. A Fourier series
We compute the Fourier series of the $1$-periodic function $\phi(t)$ from Equation (21). Let $L$ be some positive integer. Let $n\in\mathbb{Z}$. We start with
$\int_{0}^{1}\mathsf{e}^{-2\pi int}\sum_{l=-\infty}^{L}\frac{2^{t-l}}{\mathsf{e}^{2^{t-l}}-1}\mathrm{d}t=\int_{0}^{1}\mathsf{e}^{-2\pi int}\sum_{l=-L}^{\infty}\frac{2^{t+l}}{\mathsf{e}^{2^{t+l}}-1}\mathrm{d}t$
$=\int_{-L}^{+\infty}\mathsf{e}^{-2\pi int}\frac{2^{t}}{\mathsf{e}^{2^{t}}-1}\mathrm{d}t=\frac{1}{\log 2}\int_{2^{-L}}^{\infty}\mathsf{e}^{-2\pi in\log_{2}(x)}\frac{\mathrm{d}x}{\mathsf{e}^{x}-1}$
$=\frac{1}{\log 2}\int_{1}^{\infty}x^{-2\pi in(\log 2)^{-1}}\frac{\mathrm{d}x}{\mathsf{e}^{x}-1}+\frac{1}{\log 2}\int_{2^{-L}}^{1}x^{-2\pi in(\log 2)^{-1}}\Bigl(\frac{1}{\mathsf{e}^{x}-1}-\frac{1}{x}\Bigr)\mathrm{d}x$
$+\frac{1}{\log 2}\begin{cases}(-2\pi in(\log 2)^{-1})^{-1}\biggl[x^{-2\pi in(\log 2)^{-1}}\biggr]_{2^{-L}}^{1}=0&(n\neq 0),\\ L\log(2)&(n=0).\end{cases}$
Let us first handle $n=0$. We thus have:
$c_{0}(\phi)\coloneq\int_{0}^{1}\phi(t)dt$ $=\lim_{L\to\infty}\biggl(-\frac{1}{2}-\frac{1}{2}+(L+1)-\frac{1}{\log 2}\int_{2^{-L}}^{\infty}\frac{\mathrm{d}x}{\mathsf{e}^{x}-1}\biggr)$
$=\lim_{L\to\infty}\biggl(L-\frac{1}{\log 2}\bigl[\log(1-\mathsf{e}^{-x})\bigr]_{2^{-L}}^{\infty}\biggr)$
$=\lim_{L\to\infty}\log_{2}\bigl(2^{L}(1-\mathsf{e}^{-2^{-L}})\bigr)=0.$
Suppose now $n\neq 0$. We obtain
$c_{n}(\phi)$ $\coloneq\int_{0}^{1}\mathsf{e}^{-2\pi int}\phi(t)dt=\frac{1}{2\pi in}-\frac{1}{\log 2}\begin{aligned} \biggl(&\int_{1}^{\infty}x^{-2\pi in(\log 2)^{-1}}\frac{\mathrm{d}x}{\mathsf{e}^{x}-1}\\ &+\int_{0}^{1}x^{-2\pi in(\log 2)^{-1}}\Bigl(\frac{1}{\mathsf{e}^{x}-1}-\frac{1}{x}\Bigr)\mathrm{d}x\biggr)\end{aligned}$
$=\frac{1}{2\pi in}-\frac{1}{\log 2}\lim_{h\to 0^{+}}\begin{aligned} \biggl(&\int_{1}^{\infty}x^{-2\pi in(\log 2)^{-1}+h}\frac{\mathrm{d}x}{\mathsf{e}^{x}-1}\\ &+\int_{0}^{1}x^{-2\pi in(\log 2)^{-1}+h}\Bigl(\frac{1}{\mathsf{e}^{x}-1}-\frac{1}{x}\Bigr)\mathrm{d}x\biggr)\end{aligned}$
$=\frac{1}{2\pi in}-\frac{1}{\log 2}\lim_{h\to 0^{+}}\biggl(\int_{0}^{\infty}x^{-2\pi in(\log 2)^{-1}+h}\frac{\mathrm{d}x}{\mathsf{e}^{x}-1}-\int_{0}^{1}x^{-2\pi in(\log 2)^{-1}+h-1}\mathrm{d}x\biggr)$
$=\frac{1}{2\pi in}-\frac{1}{\log 2}\biggl(\Gamma\bigl(1-2\pi in(\log 2)^{-1}\bigr)\zeta\bigl(1-2\pi in(\log 2)^{-1}\bigr)-\frac{1}{-2\pi in(\log 2)^{-1}}\biggr)$
$=-(\log 2)^{-1}\Gamma\bigl(1-\frac{2\pi in}{\log 2}\bigr)\zeta\bigl(1-\frac{2\pi in}{\log 2}\bigr).$
We have used the basic Riemann formula $\Gamma(s)\zeta(s)=\int_{0}^{\infty}x^{s-1}(\mathsf{e}^{x}-1)^{-1}\mathrm{d}x$ for $\Re s>1$. Using the usual bound $|\zeta(1+it)|=O_{t\to+\infty}(\log t)$ ([12, Thm. 3.5]), and $|\Gamma(1+it)|\sim\sqrt{2\pi|t|}\exp(-\frac{\pi}{2}|t|)$ from Stirling formula in the complex domain (or more expediently from the reflection formula $\Gamma(s)\Gamma(1-s)=\pi/\sin(\pi s)$), we recover the analyticity of $\phi(t)$ in the strip $|\Im t|<\pi/\log(4)$ which was stated in Proposition 4.
This completes the proof of Theorem 2.
6. Exponential generating function
Let us point out that there is an alternative formula for $f_{m}$, $m>0$, which is more convenient numerically:
Proposition 6.
Let $f_{m}$ for real $m>0$ be defined by equation (19),i.e.
$f_{m}=\sum_{l=0}^{\infty}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr).$
There holds:
$f_{m}=\sum_{l=1}^{\infty}\frac{l2^{-l}m}{\mathsf{e}^{2^{-l}m}+1}\,.$
The author did not find the next formula in standard references [1, 7, 8] although it is undoubtly well-known, and is related to the relation
$E_{n-1}(x)=2n^{-1}\bigl(B_{n}(x)-2^{n}B_{n}(2^{-1}x)\bigr)$ (22)
between Euler and Bernoulli polynomials ([1, 23.1.27], [7, 1.14(7)]).
Lemma 2.
For $z\in\mathbb{C}\setminus 2\pi i\mathbb{Z}$,
$\sum_{k=1}^{\infty}\frac{2^{-k}z}{\exp(2^{-k}z)+1}=1-\frac{z}{\mathsf{e}^{z}-1}\,.$ (23)
Observe in passing that if we multiply both sides by $\mathsf{e}^{z}-1$ we obtain an identity of entire functions.
Proof.
For any integer $l\geq 1$ we consider the identity
$\mathsf{e}^{z}-1=(e^{z/2}+1)(\mathsf{e}^{z/4}+1)\cdots(\mathsf{e}^{z/2^{l}}+1)(\mathsf{e}^{z/2^{l}}-1),$
and compute its logarithmic derivative. This gives, for $z$ not among the poles:
$\frac{\mathsf{e}^{z}}{\mathsf{e}^{z}-1}=\sum_{k=1}^{l}\frac{\mathsf{e}^{z/2^{k}}2^{-k}}{\mathsf{e}^{z/2^{k}}+1}+\frac{\mathsf{e}^{z/2^{l}}}{2^{l}(\mathsf{e}^{z/2^{l}}-1)}.$
The limit for $l\to\infty$ gives:
$1+(\mathsf{e}^{z}-1)^{-1}=\sum_{k=1}^{\infty}\bigl(2^{-k}-\frac{2^{-k}}{\mathsf{e}^{z/2^{k}}+1}\bigr)+z^{-1},$
which, after canceling out the $1$ and multiplying by $z$ gives Equation (23). ∎
Proof of Proposition 6.
We can now compute:
$f_{m}$ $=\sum_{l=0}^{\infty}\Bigl(1-\frac{m/2^{l}}{\exp(m/2^{l})-1}\Bigr)$
$=\sum_{l=0}^{\infty}\sum_{k=1}^{\infty}\frac{2^{-l-k}m}{\exp(2^{-l-k}m)+1}$
$=\sum_{q=1}^{\infty}(\sum_{k=1}^{q}1)\frac{2^{-q}m}{\exp(2^{-q}m)+1}\,.\qed$
The statement and proof of Proposition 6 actually hold for any complex $m$ which is not in $2\pi i\mathbb{Z}$.
In the next proposition we compute (the primitive of) the exponential generating function for the $(e_{m})_{m\geq 0}$ sequence.
Proposition 7.
There holds (for $t\in\mathbb{C}$):
$\sum_{m=0}e_{m}\frac{t^{m+1}}{(m+1)!}=(\mathsf{e}^{t}-1)\sum_{l=0}^{\infty}\Bigl(1-\frac{2^{-l}t}{\mathsf{e}^{2^{-l}t}-1}\Bigr)=(\mathsf{e}^{t}-1)\sum_{l=1}^{\infty}\frac{l2^{-l}t}{\mathsf{e}^{2^{-l}t}+1}\,.$ (24)
Proof.
Note that the quantity $E(t)$ defined by the left-hand side is an entire function, and indeed if we move the factor $\mathsf{e}^{t}-1$ to inside the sums we see that the two other expressions are, too. Let us use as starting point Equation (13) (and recall $e_{0}=0$):
$\sum_{m=1}e_{m}\frac{t^{m+1}}{(m+1)!}$ $=-\sum_{m=1}\frac{t^{m+1}}{(m+1)!}\sum_{k=1}^{m}\frac{(m+1)!}{k!(m+1-k)!}B_{k}\frac{2^{k}}{2^{k}-1}$
$=-\sum_{k=1}^{\infty}\frac{B_{k}2^{k}}{k!(2^{k}-1)}\sum_{m=k}^{\infty}\frac{t^{m+1}}{(m+1-k)!}$
$=-\sum_{k=1}^{\infty}\frac{B_{k}2^{k}t^{k}}{k!(2^{k}-1)}(\mathsf{e}^{t}-1)$
$=-(\mathsf{e}^{t}-1)\sum_{k=1}^{\infty}B_{k}\frac{t^{k}}{k!}\sum_{l=0}^{\infty}2^{-lp}$
$=-(\mathsf{e}^{t}-1)\sum_{l=0}^{\infty}\Bigl(\frac{2^{-l}t}{\mathsf{e}^{2^{-l}t}-1}-1\Bigr).$
This gives the first equality from Equation (24). The second one was already stated, in another variable, as Proposition 19. ∎
Remark 6.
Hence, if we let $E(t)$ denote the left-hand side of Equation (24), we see that $E(m)=(\mathsf{e}^{m}-1)f_{m}=(\mathsf{e}^{m}-1)\bigl(e_{m}+O(\frac{\log m}{m})\bigr)=\mathsf{e}^{m}e_{m}\bigl(1+O(m^{-1})\bigr)$ and thus $e_{m}=(1+O(m^{-1}))\mathsf{e}^{-m}E(m)=\mathsf{e}^{-m}E(m)+O(\frac{\log m}{m})$. This is very reminiscent of the general estimate from [6, Prop. 1] relating moments of some measures on the unit interval with their exponential generating function.
Remark 7.
A form of the functional equation verified by $E(t)$ (which has neither constant term nor a linear term, as $e_{0}=0$) is
$E(2t)=(\mathsf{e}^{t}+1){E(t)}+\mathsf{e}^{2t}-1-2t.$ (25)
Looking at (25) as an identity of formal power series, it implies that $E$ has no constant term, but not that $E^{\prime}(0)=0$. Equation (25) with the definition $E(t)=\sum_{m=0}^{\infty}e_{m}\frac{t^{m+1}}{(m+1)!}$ is equivalent to the recurrence relation Equation (1) but stated without the initial condition $e_{0}=0$.
7. Concluding remarks
The $\phi$ function from Theorem 2 fits into an old theme going back to Hardy [9] (see also [10]). Balazard et al. [3] have studied under some general conditions series of the type $f_{\alpha}(x)=\sum_{n\in\mathbb{Z}}\alpha(x^{\theta^{n}})$ where $\alpha$ is continuous on $[0,1]$, vanishes at 0 and $1$ and has a power series expansion $\alpha(x)=\sum_{k=1}^{\infty}a_{k}x^{k}$ (convergent for $0\leq x<1$). They obtain the Fourier series of $f_{\alpha}$ as a function of $\log_{\theta}\log x^{-1}$ in terms of values at $(\log\theta)^{-1}2\pi i\mathbb{Z}\setminus\{0\}$ of the Gamma function and a Dirichlet series associated to $\alpha$. This is reminiscent of the computations done in section 5 (we did not need to use the Poisson summation formula).
They consider in particular the case $\alpha(x)=x\prod_{n=0}^{\infty}(1-x^{2^{n}})$, which is related to the Thue-Morse sequence (see also about this [2]). It is of note that the function $\phi(t)$ ($t=\log_{2}(m)$) from Theorem 2 and Equation (21) is related to the logarithmic derivative of $m\mapsto\mathsf{e}^{m}\alpha(\mathsf{e}^{-m})$. Indeed, we compute straightforwardly:
$m\frac{\mathrm{d}}{\mathrm{d}m}\log\prod_{n=0}^{\infty}(1-\mathsf{e}^{-2^{n}m})=m\sum_{n=0}^{\infty}\frac{2^{n}\mathsf{e}^{-2^{n}m}}{1-e^{-2^{n}m}}=\sum_{n=0}^{\infty}\frac{2^{n}m}{\mathsf{e}^{2^{n}m}-1}=\frac{m}{\mathsf{e}^{m}-1}-g_{m}\,.$
Here $g_{m}$ is the quantity defined in Equation (20).
This computation motivates the consideration of the following infinite product:
$\Psi(z)=2^{\frac{\operatorname{Log}_{2}(z)(\operatorname{Log}_{2}(z)+1)}{2}}\prod_{l=0}^{\infty}\frac{1-\mathsf{e}^{-2^{-l}z}}{2^{-l}z}\prod_{k=1}^{\infty}(1-\mathsf{e}^{-2^{k}z}).$ (26)
The first infinite product is an entire function. The second one is analytic in the half-plane $\Re z>0$. So $\Psi$ is analytic there. It verifies the functional equation:
$\Psi(2z)=2^{\frac{(\operatorname{Log}_{2}(z)+1)(\operatorname{Log}_{2}(z)+2)}{2}-\frac{\operatorname{Log}_{2}(z)(\operatorname{Log}_{2}(z)+1)}{2}}\Psi(z)\frac{(1-\mathsf{e}^{-2z})/(2z)}{1-\mathsf{e}^{-2z}}=\Psi(z)$
A straightforward computation gives the logarithmic derivative with respect to the variable $\operatorname{Log}z$:
$z\frac{\mathrm{d}}{\mathrm{d}z}\log\Psi(z)=\operatorname{Log}_{2}(z)+\frac{1}{2}+\sum_{l=0}^{\infty}\Bigl(\frac{2^{-l}z}{\mathsf{e}^{2^{-l}z}-1}-1\Bigr)+\sum_{k=1}^{\infty}\frac{2^{k}z}{\mathsf{e}^{2^{k}z}-1}.$
Comparison with Equation (21) shows that the above is exactly $\phi(\operatorname{Log}_{2}z)$. And indeed, the image of the right half-plane under $z\mapsto\operatorname{Log}_{2}z$ is the horizontal strip where we have defined $\phi$ as an nalytic function.
Whether this relation between $\Psi$ and $\phi$ is indicative of some closer connection between the sequence $(e_{m})$ and the Thue-Morse sequence [2, 3] is currently unknown to the author.

### 15 May 2026 v4

@@
 Key words and phrases:
-Euler’s constant, Euler-Mascheroni constant, asymptotic of sequences
+Euler’s constant, asymptotics, Riemann zeta function, Bernoulli numbers and polynomials, linear recurrences involving binomial coefficients
 2020 Mathematics Subject Classification:
-11Y60, 11B83 (Primary) 11B68, 11M41, 40G99 (Secondary)
+11Y60, 11B83 (Primary) 11B68, 11M41, 40G99, 60C05 (Secondary)
 1. Main result
@@
 $e_{m}=\frac{2^{m+1}+\sum_{j=1}^{m}\binom{m+1}{j}e_{m-j}}{2^{m+1}-2}.$ (1)
-This gives the sequence $2$, $\frac{7}{3}$, $\frac{8}{3}$, $\frac{133}{45}$, $\frac{16}{5}$, … , $e_{10}=\frac{163287}{40579}$, …. See Table 1 for more values.
-Table 1. The coefficients $e_{m}$, $1\leq m\leq 20$
+This gives the sequence 0, 22, $\frac{7}{3}$, $\frac{8}{3}$, $\frac{133}{45}$, $\frac{16}{5}$, … , $e_{10}=\frac{163287}{40579}$, …. See Table 1 for more values.
+Table 1. The coefficients $e_{m}$, $1\leq m\leq 20$.
+See also https://oeis.org/A372422.
 $m$ $e_{m}$
@@
 2020 $\frac{23101294621895391907711}{4689192129680103420375}$
-As usual, $\gamma$ is Euler constant. Here is the main Theorem.
+They can be expressed, as we will see, in terms of Bernoulli numbers (using the convention $B_{1}=-\frac{1}{2}$):
+$e_{m}=-\sum_{k=1}^{m}\binom{m+1}{k}B_{k}\frac{2^{k}}{2^{k}-1}=1-\sum_{k=1}^{m}\binom{m+1}{k}B_{k}\frac{1}{2^{k}-1}\,.$ (2)
+(the second formulation supposes $m\geq 1$).
+Let $\gamma$ be as usual the Euler-Mascheroni constant. Here is the main Theorem:
 Theorem 1.
-Let $\ell\geq 2$. Let $(e_{m})$ be the sequence defined by (1). There holds
+Let $\ell\geq 2$. Let $(e_{m})$ be the sequence defined by Equation (1). There holds
 $\gamma=\sum_{n=1}^{2^{\ell-1}-1}\frac{1}{n}-(\ell-1)\log 2+\sum_{m=1}^{\infty}\frac{(-1)^{m-1}e_{m}}{m+1}\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{m+1}}.$ (3)
@@
 The level $\ell$ is the number of binary digits used by the integers whose inverse powers are added in Equation (3).
-In Table 2, we illustrate the numerical values obtained using up to $e_{10}$ for $\ell$ being equal to $2$, 33, …, $7$ (so that the computations could have been done by a diligent human calculator), and also up to $e_{20}$ for $\ell=2,3,4$.
-The computed approximations to Euler’s constant $\gamma$ in this table are truncated not rounded.
-Table 2. Table of approximations of $\gamma$
+In Table 2, we illustrate the result of computations well within the reach of a diligent human calculator: they use each $10$ terms of the series for $\ell$ respectively equal to $2$, 33, …, $7$, and $20$ terms for $\ell=2,3,4$.
+Table 2. Table of approximations of $\gamma$ (decimal expansions in the middle column are truncated)
 $\ell$ Series from Thm. 1 up to $e_{m}=e_{10}$ last term
@@
 A numerical implementation using Python and the mpmath library is provided at https://burnolmath.gitlab.io/dyadic-gamma/ (it uses by default $\ell=8$).
-One can express (see Equation (13)) $e_{m}$ as an explicit finite linear combination of the Bernoulli numbers $B_{k}$, $1\leq k\leq m$, but this uses terms which are much larger than $e_{m}$ itself.
-So, it seems that using the recurrence (1), which is stable numerically, is more appropriate for computing $e_{m}$’s with hundreds of decimal places.
-However, an algorithm using (1) appears to have an intrinsic (at least) quadratic cost, and is not suited for large scale computations.
-[…]
+The coefficients $e_{m}$ are computed using the recurrence (1) (the larger the $m$, the less the required precision), which induces a quadratic cost which is redhibitory for large scale computations.
+Equation (2) is not immediately usable numerically for large $m$, as individual terms are much larger than $e_{m}$.
+Indeed it turns out that $e_{m}\sim\log_{2}(m)$, more precisely the difference is bounded.
+See Table 3 and Figure 1, which displays intriguing oscillations of $e_{m}-H_{m+1}/\log(2)$ ($H_{n}=\sum_{1\leq j\leq n}j^{-1}$) as a function of $\log_{2}(m)$.
 Table 3. Comparison of $e_{m}$ with $(\log 2)^{-1}H_{m+1}$
@@
 2020 -0.332 643 144 549
-Figure 1. $e_{m}-(\log 2)^{-1}H_{m+1}$ vs $\log_{2}(m)$, $16\leq m\leq 500$
+Figure 1. $e_{m}-(\log 2)^{-1}H_{m+1}$ as function of $\log_{2}(m)$, $16\leq m\leq 500$. Notice the oscillations as function of $\log_{2}(m)$ and their small amplitudes.
+This is the time to reveal that the sequence $(e_{m})$ (and its peculiar asymptotic) has long been known in the field of theoretical computer science!
+The numerators and denominators have dedicated OEIS pages A372422 and A372423.
+Paraphrasing their description in a less precise way, $e_{m}$ is the expected depth of trees related to the process of recursively randomly eliminating people via coin tosses, from a group of initial cardinality $m+1$.
+It is one among a family of related sequences arising in the theoretical analysis of search algorithms, to which Knuth dedicated an entire chapter of his influential treatise ([14, §6.3 “Digital searching”]).
+See also the article [9] by Flajolet and Sedgewick which contains a discussion of certain techniques for asymptotic analysis involving contour integrals.
+We were led to the sequence $(e_{m})$ via the analysis of zeta series with missing digits (inclusive of the case with no missing digits…) [4, 5], so the occurrence in the context of “digital” search algorithm is in retrospect perhaps not so surprising.
+The quantity $e_{m}$ is $d_{m+1}$ from [16, §4].
+In this reference, Prodinger establishes an asymptotic $d_{m}\sim\log_{2}(m)+\frac{1}{2}-\delta_{2}(\log_{2}(m))$ ([16, Thm.
+14]) where $\delta_{2}$ is a $1$-periodic function of zero average and small amplitude, given as a Fourier series involving the values of $\Gamma(s)\zeta(s)$ on the line $\Re(s)=1$.
+The order of the error is not indicated in [16].
+Earlier, Szpankowski published a general theorem about binomial sums with Bernoulli coefficients of the type of Equation (2).
+This Theorem 4.1 of [17] claims a general error bound which would translate here in our notation into $O(m^{-1})$.
+Unfortunately, this proof has a gap.
+[^1] In the study of $\delta_{1}(n,s)$ of [17, eq.
+(4.12)], the estimate $\binom{n+1}{k}n^{-k}=k!^{-1}\bigl(1+O(n^{-1})\bigr)$ is used for $1\leq k\leq n$ but is of course valid only for bounded $k$’s, or at least under some restriction on $k$ relative to $n$.
+Hence, the proof given for the total contribution for $j>\log_{d}(n)$ is invalid.
+We shall explain, in our context, how to fix it.
+Knuth had studied earlier in his treatise [14] similar such combinations of Bernoulli numbers.
+A particular example is the topic of [14, Ex.
+6.3-34] (which is not about the sequence $e_{m}$, but a related one).
+There, an asymptotic formula is obtained with a claimed $O(n^{-1})$ bound for the error.
+But the proof given ([14, page 727]) is not valid.
+[^2] The exercise [14, 6.4-34, p.
+510] (page 503 in the first edition) tells the reader to prove that a series $\sum_{j}\bigl(S_{n-1}(2^{j})-T_{n}(2^{j})\bigr)$ is $O(n^{-1})$ (using here notation from the answer).
+The answer on page 727 however appears to be about $\sum_{j}\bigl(S_{n}(2^{j})-T_{n}(2^{j})\bigr)$.
+But this latter series badly diverges, its general term goes to -$\infty$!
+Besides, the claim $S_{n}(m)=T_{n}(m)+O(\mathsf{e}^{-n/m}n/m^{2})$ is wrong (it is correct if $m/n$ is bounded).
+The $O(n^{-1})$ bound itself though is correct, as we shall establish.
+We shall give a proof for this $O(n^{-1})$.
+We give a detailed self-contained approach for the $(e_{m})$ asymptotic.
 Theorem 2.
-There holds
-$e_{m}=\log_{2}(m)+\frac{1}{2}+\phi(\log_{2}(m))+O(\frac{\log m}{m})$ (3)
-with
-$\phi(t)=-(\log 2)^{-1}\sum_{n\in\mathbb{Z},n\neq 0}\Gamma\bigl(1-\frac{2\pi in}{\log 2}\bigr)\zeta\bigl(1-\frac{2\pi in}{\log 2}\bigr)\mathsf{e}^{2\pi int}.$
-[…]
-Figure 2. $e_{m}-(\log 2)^{-1}H_{m+1}$ as function of $\log_{2}(m)$, and the graph of $\frac{\gamma}{\log 2}+\frac{1}{2}+\phi(t)$, $16\leq m\leq 1000$
-As a corollary to the Theorem, $e_{2m}\approx 1+e_{m}$, a fact which does not seem immediately obvious from Equation 1.
+The sequence $(e_{m})$ obeys the following asymptotic:
+$e_{m}=\log_{2}(m)+\frac{1}{2}+\phi(\log_{2}(m))+O(m^{-1}),$ (4)
+where $\phi$ is a $1$-periodic function, which is analytic for $|\Im t|<\pi/(2\log(2)$), has zero mean, and verifies
+$\phi(t)=-t-\frac{1}{2}+\sum_{l=0}^{\infty}\bigl(1-\frac{2^{t-l}}{\mathsf{e}^{2^{t-l}}-1}\bigr)-\sum_{l=-1}^{-\infty}\frac{2^{t-l}}{\mathsf{e}^{2^{t-l}}-1}\,$, (5)
+and
+$\phi(t)=-(\log 2)^{-1}\sum_{n\in\mathbb{Z},n\neq 0}\Gamma\bigl(1-\frac{2\pi in}{\log 2}\bigr)\zeta\bigl(1-\frac{2\pi in}{\log 2}\bigr)\mathsf{e}^{2\pi int}\,.$ (6)
+Figure 2. $e_{m}-(\log 2)^{-1}H_{m+1}$ as function of $\log_{2}(m)$, and the graph of $\frac{\gamma}{\log 2}+\frac{1}{2}+\phi(t)$, $16\leq m\leq 1000$, $t=\log_{2}(m)$.
 Remark 1.
-Numerical investigations suggest that the logarithm in the numerator of the error term from Equation (3) can be omitted.
-However we should remain cautious because we can compute in reasonable time only of the order of ten thousand coefficients.
-To be more precise, the error term appears to have the shape $\psi(\log_{2}(m))/m+o(1/m)$ with some $1$-periodic function $\psi$ averaging to $\frac{3}{2\log(2)}$.
+Due to the $O(m^{-1})$ error term it does not matter here whether we are talking about $e_{m}$ ($d_{m+1}$) or about $d_{m}$ from [16, Thm. 14].
+Remark 2.
+Numerical computations suggest that the $O(m^{-1})$ error term in (4) is $m^{-1}\psi(\log_{2}(m))+o(m^{-1})$ with some $1$-periodic function $\psi$ averaging to $\frac{3}{2\log(2)}$ (whether we use $m$ or $m+1$ now matters for what $\psi$ actually is).
+Related to this, for $m$ up to a few thousands, $\frac{H_{m+1}}{\log(2)}-\frac{\gamma}{\log 2}+\frac{1}{2}$, which differs from $\log_{2}(m)+\frac{1}{2}$ by a term equivalent to $\frac{3}{2\log(2)m}$, proves to be a much better approximation to $e_{m}$ than $\log_{2}(m)+\frac{1}{2}$.
+Remark 3.
+We shall obtain the Fourier series (6) as a corollary to (5), which itself will be derived by elementary means.
+This is thus quite different from the contour integral approach sketched by Prodinger in [16] and Szpankowski [17].
+On discovering the gaps regarding error bounds, we expected that the methods of complex analysis, for which an account is to be found in [9] by Flajolet and Sedgewick (see also [11] and the book [10] for many resources), would be the tools of choice.
+As it turns out, we obtain here Equation (4) with its $O(m^{-1})$ error bound with no recourse to them.
+It is probable though that they could prove of use for refinements.
 Remark 4.
 […]
-Remark 3.
-We will establish concrete explicit upper and lower bounds for $e_{m}$. They are useful, especially the upper bound, to calibrate numerical implementations.
 2. Geometric series for $\gamma$
@@
 $c_{m}(s)=\frac{1}{2^{m+s}-2}\sum_{j=1}^{m}\binom{m}{j}c_{m-j}(s).$ (7)
-In [5], there is an integer parameter $b>1$, which here has been set to $b=2$.
-There are generalizations of Theorem 1 to general $b>2$, where the analog of the recurrence from Equation (1) has additional factors given by the power sums of the base-$b$ digits.
+In [5], there is more generally an integer parameter $b>1$, (which would be $M$ in the notation of [14, §6.3]) which in Equation (7) has been set to $b=2$.
+And there are generalizations of Theorem 1 to general $b>2$, where the analog of the recurrence from Equation (1) has additional factors given by the power sums of the base-$b$ digits.
+Already in the Treatise of Knuth [14, §6.3], such equations arise, with an integer $M>1$ like the $b$ of [5].
 Let $\ell\geq 2$, the level, be some integer greater than one.
@@
 This, indeed, is Equation (1). And $e_{0}=0$. ∎
-3. Estimates for the coefficients $e_{m}$
-In this section, we establish the next proposition as a corollary to the recurrence from Equation (1).
+3. Elementary bounds
+In this section, we mention two elementary estimates which can be established using only the recurrence (1).
+As such results are superseded for large $m$ (if one goes through the proof to get explicit bounds) by Theorem 2 and by explicit numerical results for $m$ of moderate size, we omit the proofs and refer the interested reader to https://arxiv.org/abs/2603.29998v1.
 Proposition 1.
-For $m\geq 2$, there holds
-$\frac{H_{m+1}}{\log 2}-\numprint{0.35}<e_{m}<\frac{H_{m+1}}{\log 2}-\numprint{0.31}\;.$
-We first establish a less tight bound. We will then re-use the tools from the proof of the Lemma 1 to prove Proposition 1.
-Lemma 1.
 There holds for any $m\in\mathbb{N}$ (with $H_{n}=\sum_{j=1}^{n}j^{-1}$):
-$\frac{H_{m+1}-1}{\log 2}\leq e_{m}<\frac{H_{m+1}}{\log 2}-\numprint{0.161}\,.$
-Proof.
-Let us define $\delta_{n}=e_{n}-(\log 2)^{-1}H_{n+1}$ for $n\geq 0$. We thus have, for $m\geq 1$:
-$(2^{m+1}-2)e_{m}=\left\{\begin{aligned} 2^{m+1}&+(\log 2)^{-1}\sum_{j=1}^{m}\binom{m+1}{j}H_{m-j+1}\\ &+\sum_{j=1}^{m}\binom{m+1}{j}\delta_{m-j}.\end{aligned}\right$. (8)
-Now (as is well-known):
-$H_{m+1}+\sum_{j=1}^{m}\binom{m+1}{j}H_{m-j+1}$ $=\sum_{k=1}^{m+1}\binom{m+1}{k}H_{k}$
-$=\int_{0}^{1}\sum_{k=0}^{m+1}\binom{m+1}{k}\frac{1-t^{k}}{1-t}\mathrm{d}t$
-$=\int_{0}^{1}\frac{2^{m+1}-(1+t)^{m+1}}{1-t}\mathrm{d}t$
-$=\int_{1}^{2}\frac{2^{m+1}-u^{m+1}}{2-u}\mathrm{d}u$
-$=2^{m+1}\int_{\frac{1}{2}}^{1}\frac{1-x^{m+1}}{1-x}\mathrm{d}x$
-$=2^{m+1}H_{m+1}-2^{m+1}\int_{0}^{\frac{1}{2}}\frac{1-x^{m+1}}{1-x}\mathrm{d}x$
-$=2^{m+1}H_{m+1}-2^{m+1}\log 2+2^{m+1}\underbrace{\int_{0}^{\frac{1}{2}}\frac{x^{m+1}}{1-x}\mathrm{d}x}_{\epsilon_{m}}.$
-Hence, from (8),
-$(2^{m+1}-2)e_{m}=(2^{m+1}-2)\frac{H_{m+1}}{\log 2}+\frac{H_{m+1}}{\log 2}+2^{m+1}\frac{\epsilon_{m}}{\log 2}+\sum_{j=1}^{m}\binom{m+1}{j}\delta_{m-j}.$
-$e_{m}=\frac{H_{m+1}}{\log 2}+\underbrace{\frac{H_{m+1}+2^{m+1}\epsilon_{m}}{(2^{m+1}-2)\log 2}+\frac{\sum_{j=1}^{m}\binom{m+1}{j}\delta_{m-j}}{2^{m+1}-2}}_{\delta_{m}}$ (9)
-Suppose now by induction that $\delta_{j}\geq-(\log 2)^{-1}$, for $0\leq j<m$, which is true for $m=1$, as $\delta_{0}=-(\log 2)^{-1}$.
-The above implies $\delta_{m}\geq-(\log 2)^{-1}$ too, so it is true for all $m$.
-To obtain an upper bound, we define
-$\mu_{m}=\max_{0\leq j\leq m}\delta_{j}.$
-In particular $\mu_{0}=\delta_{0}=-(\log 2)^{-1}\approx\numprint{-1.443}$, and $\mu_{1}=\delta_{1}=2-(\log 2)^{-1}\frac{3}{2}\approx\numprint{-0.164}$. The above implies, for $m\geq 1$:
-$\mu_{m}\leq\frac{H_{m+1}+2^{m+1}\epsilon_{m}}{(2^{m+1}-2)\log 2}+\mu_{m-1},$
-hence by induction, we obtain, for any given $m$:
-$p>m$ $\implies\mu_{p}\leq\mu_{m}+(\log 2)^{-1}\sum_{n=m+1}^{p}\frac{H_{n+1}+2^{n+1}\epsilon_{n}}{2^{n+1}-2},$
-thus, using
-$\frac{1}{2^{n+1}-2}$ $=\frac{1}{2^{n+1}}+\frac{2}{2^{n+1}(2^{n+1}-2)}\leq\frac{1}{2^{n+1}}+\frac{1}{4^{n}},$
-we obtain, for any $p>m\geq 0$:
-$\delta_{p}\leq\mu_{m}+(\log 2)^{-1}\sum_{n=m+1}^{\infty}\left(2^{-n-1}H_{n+1}+4^{-n}H_{n+1}+\epsilon_{n}+2^{-n+1}\epsilon_{n}\right).$
-Now,
-$\sum_{n=m+1}^{\infty}2^{-n-1}H_{n+1}=\frac{-\log(1-\frac{1}{2})}{1-\frac{1}{2}}-\sum_{n=0}^{m}2^{-n-1}H_{n+1},$
-and
-$\sum_{n=m+1}^{\infty}4^{-n}H_{n+1}=4\frac{-\log(1-\frac{1}{4})}{1-\frac{1}{4}}-\sum_{n=0}^{m}4^{-n}H_{n+1}.$
-Evaluating numerically for $m=10$ we obtain respectively:
-$(\log 2)^{-1}\sum_{n=11}^{\infty}2^{-n-1}H_{n+1}$ $\approx\numprint{0.002237}$
-[…]
-We bound trivially $\epsilon_{n}=\int_{0}^{\frac{1}{2}}x^{n+1}(1-x)^{-1}\mathrm{d}x<2^{-n-1}$, hence:
-$(\log 2)^{-1}\sum_{n=11}^{\infty}(1+2^{-n+1})\epsilon_{n}$ $<(\log 2)^{-1}(1+2^{-10})\sum_{n=11}^{\infty}2^{-n-1}$
-$=(\log 2)^{-1}(2^{-11}+2^{-21})\approx\numprint{0.000705}$
-So for $p>10$, $\delta_{p}<\mu_{10}+\numprint{0.003}$.
-A quick look at Table 3 shows that $\mu_{10}=\delta_{1}=2-(\log 2)^{-1}\frac{3}{2}\approx\numprint{-0.164043}$.
-Hence $\mu_{10}+\numprint{0.003}<\numprint{-0.161}$.
-This concludes the proof of Lemma 1.
-∎
-Proof of Proposition 1.
-We continue with the same technique as in the previous proof. To improve the upper bound, we separate out $\delta_{1}$:
-$(2^{m+1}-2)^{-1}\sum_{j=1}^{m}\binom{m+1}{j}\delta_{m-j}$ $\leq\frac{m(m+1)}{2(2^{m+1}-2)}\delta_{1}$
-$\mkern-100.0mu+\frac{2^{m+1}-2-m(m+1)/2}{2^{m+1}-2}\max\{\delta_{j},0\leq j<m,j\neq 1\}.$
-As we know that $\delta_{j}\geq\delta_{0}=-(\log 2)^{-1}$, we can further bound this as
-$\leq\frac{m(m+1)}{2(2^{m+1}-2)}\bigl(\underbrace{\delta_{1}+|\delta_{0}|}_{2-\frac{1}{2}(\log 2)^{-1}\approx\numprint{1.279}}\bigr)+\max\{\delta_{j},0\leq j<m,j\neq 1\}.$
-We again use $(2^{n+1}-2)^{-1}<2^{-n-1}(1+2^{-n+1})$ for $n\geq 1$. So:
-$\sum_{n=21}^{\infty}\frac{n(n+1)}{2(2^{n+1}-2)}$ $<(1+2^{-20})\frac{1}{8}\sum_{n=21}^{\infty}\frac{n(n+1)}{2^{n-1}}$
-$=(1+2^{-20})\frac{1}{8}\left(\frac{2}{(1-\frac{1}{2})^{3}}-\sum_{k=0}^{20}\frac{k(k+1)}{2^{k-1}}\right)$
-$\approx\numprint{0.000121117}$
-The product of this with $\delta_{1}-\delta_{0}$ is less than $\numprint{0.0002}$.
-We still need to take into account the contribution to the upper bound from the series with the $H_{n+1}$’s and the $\epsilon_{n}$’s.
-We expect something about one thousand times smaller than what we computed in the previous proof for the remainders, and indeed, one obtains:
-[…]
-[…]
-$(\log 2)^{-1}\sum_{n=21}^{\infty}(1+2^{-n+1})\epsilon_{n}$ $<(\log 2)^{-1}(1+2^{-20})\sum_{n=21}^{\infty}2^{-n-1}$
-[…]
-As $\delta_{2}\approx\numprint{-0.311608}$ is, as we see from Table 3, the maximum among the $\delta_{m}$’s, for $m\neq 1$ and $0\leq m\leq 20$, we conclude, with some margin, that
-$m\geq 2\implies\delta_{m}<\numprint{-0.3114}\;.$
-Similarly, regarding the lower bound, we can separate out $\delta_{0}=-(\log 2)^{-1}$, writing
-$(2^{p+1}-2)^{-1}\sum_{j=1}^{p}\binom{p+1}{j}\delta_{p-j}\geq\frac{p+1}{(2^{p+1}-2)}\delta_{0}+\min\{\delta_{j},1\leq j<p\},$
-where we used the fact that we know $\delta_{j}\leq 0$ for all $j$’s.
-Now working again with $p>10$, we obtain (as the other contributions involving the $H_{n+1}$’s and the $\epsilon_{n}$’s are positive):
-$p>10\implies\delta_{p}\geq\delta_{0}\sum_{n=11}^{\infty}\frac{n+1}{(2^{n+1}-2)}+\underbrace{\min\{\delta_{j},1\leq j\leq 10\}}_{{}=\delta_{3}}.$
-Now
-$\sum_{n=11}^{\infty}\frac{n+1}{(2^{n+1}-2)}$ $<(1+2^{-10})\sum_{n=11}^{\infty}\frac{n+1}{2^{n+1}}$
-$=(1+2^{-10})\frac{1}{2}\left(\frac{1}{(1-\frac{1}{2})^{2}}-\sum_{k=0}^{10}(k+1)2^{-k}\right)$
-$\approx\numprint{0.006354}.$
-The product of this with $\delta_{0}$ is about $\numprint{-0.009167}>\numprint{-0.01}$ and we thus can conclude that for all $m\geq 1$, $\delta_{m}>\numprint{-0.01}+\delta_{3}>\numprint{-0.35}$. ∎
-4. The asymptotic behavior of $(e_{m})$
+$\frac{H_{m+1}-1}{\log 2}\leq e_{m}<\frac{H_{m+1}}{\log 2}-\numprint{0.161}\,.$ (11)
+More precisely, for $m\geq 2$:
+$\frac{H_{m+1}}{\log 2}-\numprint{0.35}<e_{m}<\frac{H_{m+1}}{\log 2}-\numprint{0.31}\;.$ (12)
+4. The coefficients $e_{m}$ via power sums
 We have seen in the proof of Theorem 1 that we can define $e_{m}$ as $(m+1)c_{m}^{\prime}(1)/\log 2$ where the meromorphic functions $c_{m}(s)$ verify the recurrence Equation (7).
 From [5] and [4], there holds $c_{m}(s)=\frac{2^{s}-2}{2^{s}}u_{m}(s)$, where the meromorphic function $u_{m}(s)$ is, for $\Re s>1$, the $m$th moment of a certain discrete (complex) measure on $[0,1]$:
-$\Re s>1\implies u_{m}(s)=\sum_{l=0}^{\infty}\biggl(\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)2^{-ls}\,.$ (10)
+$\Re s>1\implies u_{m}(s)=\sum_{l=0}^{\infty}\biggl(\sum_{0\leq k<2^{l}}\bigl(\frac{k}{2^{l}}\bigr)^{m}\biggr)2^{-ls}\,.$ (13)
 In particular $u_{0}(s)=\sum_{l=0}^{\infty}2^{-l(s-1)}=2^{s}/(2^{s}-2)$. As the function $t\mapsto t^{m}$ is non-decreasing and convex, there holds
-$0\leq\int_{0}^{1}t^{m}\mathrm{d}t-2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\leq\frac{1}{2}2^{-l}.$ (11)
-So, at first for $\Re s=1$:
-$u_{m}(s)$ $=(m+1)^{-1}\sum_{l=0}^{\infty}2^{-l(s-1)}-\sum_{l=0}^{\infty}\biggl((m+1)^{-1}-2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}$
-$(m+1)u_{m}(s)$ $=\frac{2^{s}}{2^{s}-2}-\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}$
-$(m+1)c_{m}(s)$ $=1-\frac{2^{s}-2}{2^{s}}\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}.$
-Thanks to estimate (11), this gives the analytic continution to $\Re s>0$, and we can now compute the value of the derivative at $s=1$:
-$(m+1)c_{m}^{\prime}(1)=-\log(2)\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr).$
+$0\leq\int_{0}^{1}t^{m}\mathrm{d}t-2^{-l}\sum_{0\leq k<2^{l}}\bigl(\frac{k}{2^{l}}\bigr)^{m}\leq\frac{1}{2}2^{-l}.$ (14)
+So, at first for $\Re s>1$:
+$u_{m}(s)$ $=(m+1)^{-1}\sum_{l=0}^{\infty}2^{-l(s-1)}-\sum_{l=0}^{\infty}\biggl((m+1)^{-1}-2^{-l}\sum_{0\leq k<2^{l}}\bigl(\frac{k}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}$
+$(m+1)u_{m}(s)$ $=\frac{2^{s}}{2^{s}-2}-\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq k<2^{l}}\bigl(\frac{k}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}$
+$(m+1)c_{m}(s)$ $=1-\frac{2^{s}-2}{2^{s}}\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq k<2^{l}}\bigl(\frac{k}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}.$
+Thanks to estimate (14), this gives the analytic continuation to $\Re s>0$, and we can now compute the value of the derivative at $s=1$:
+$(m+1)c_{m}^{\prime}(1)=-\log(2)\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq k<2^{l}}\bigl(\frac{k}{2^{l}}\bigr)^{m}\biggr).$
 From Equation (10) which arose in the course of the proof of Theorem 1, we thus get an “explicit” formula for the coefficients $e_{m}$:
@@
 There holds:
-$e_{m}=\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)\,.$ (12)
+$e_{m}=\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq k<2^{l}}\bigl(\frac{k}{2^{l}}\bigr)^{m}\biggr)\,.$ (15)
 In terms of Bernoulli polynomials and numbers:
 $e_{m}=\sum_{l=0}^{\infty}\biggl(1-\frac{B_{m+1}(2^{l})-B_{m+1}(0)}{2^{(m+1)l}}\biggr)=-\sum_{k=1}^{m}\binom{m+1}{k}B_{k}\frac{2^{k}}{2^{k}-1}\,.$ (16)
+To obtain Equation (16), we use the properties of Bernoulli numbers and polynomials as summarized below, and $2^{k}/(2^{k}-1)$ is obtained as $\sum_{l=0}^{\infty}2^{-kl}$.
+As $B_{m+1}(1)-B_{m+1}(0)=0$ if $m>0$, the $l=0$ term contributes +1+1 to $e_{m}$, and there also holds:
+$m\geq 1\implies e_{m}=1-\sum_{k=1}^{m}\binom{m+1}{k}B_{k}\frac{1}{2^{k}-1}\,.$
+The Bernoulli polynomials and numbers used in (16) are defined by
+$\frac{t\mathsf{e}^{xt}}{\mathsf{e}^{t}-1}=\sum_{n=0}^{\infty}B_{n}(x)\frac{t^{n}}{n!}.$
+They verify $B_{n}(x+1)-B_{n}(x)=nx^{n-1}$, hence
+$\sum_{0\leq k<n}k^{m}=\frac{B_{m+1}(n)-B_{m+1}(0)}{m+1}\,.$
+Also
+$B_{m+1}(x)=\sum_{k=0}^{m+1}\binom{m+1}{k}B_{k}x^{m+1-k}\,$,
+and $B_{0}(x)=1$, $B_{1}(x)=x-\frac{1}{2}$, $B_{n}=B_{n}(0)$, $B_{2p}=(-1)^{p-1}|B_{2p}|$ ($p\geq 1$), $B_{2p+1}=0$ ($p\geq 1$).
+That $B_{n}$ is used both for the polynomial and its value at zero should not cause confusion.
 Remark 5.
-The Bernoulli polynomials and numbers used in (13) are defined by $\sum_{n=0}^{\infty}B_{n}(x)\frac{t^{n}}{n!}=\frac{t\mathsf{e}^{xt}}{\mathsf{e}^{t}-1}$, $B_{n}(x)=\sum_{k=0}^{n}\binom{n}{k}B_{k}x^{n-k}$, $B_{0}=1$, $B_{1}=-\frac{1}{2}$, $B_{n}=B_{n}(0)$, $(m+1)\sum_{0\leq n<N}n^{m}=B_{m+1}(N)-B_{m+1}(0)$.
-Formula (13) has large contributions of both signs and does not look promising numerically.
-In the sequel we will only use (12).
+A variant of (16) is:
+$e_{m}=\sum_{l=0}^{\infty}(l+1)\frac{B_{m+1}(2^{l+1})-2^{m+1}B_{m+1}(2^{l})+(2^{m+1}-1)B_{m+1}(0)}{2^{(l+1)(m+1)}}$ (17)
+One only needs to write the numerator as $(l+1)(v_{l+1}-2^{m+1}v_{l})$ with $v_{l}=B_{m+1}(2^{l})-B_{m+1}(0)-2^{(m+1)l}$ and rearrange the sum.
+Details are left to the reader.
+In terms of the Euler polynomials (see Equation (32)), this is:
+$\frac{m+1}{2}\sum_{l=0}^{\infty}(l+1)\frac{E_{m}(2^{l+1})-E_{m}(0)}{2^{(l+1)(m+1)}}\,.$
+5. Approximations
+Let us make the general definition, for $n$ a positive integer:
+$p_{m}(n)=n^{-1}\sum_{0\leq k<n}(n^{-1}k)^{m}\,.$ (18)
+They are the moments of the discrete probability measure $\sum_{0\leq k<n}n^{-1}\delta_{k/n}$.
+It is not needed here for $m$ to be an integer, and in what follows we let $e_{m}$ for $m$ real positive be given by Equation (15), i.e.:
+$e_{m}=\sum_{l=0}^{\infty}\bigl(1-(m+1)p_{m}(2^{l})\bigr)\,$, (19)
+and we study the asymptotic for real positive $m$ going to infinity.
+There holds
+$(m+1)p_{m}(n)=\frac{B_{m+1}(n)-B_{m+1}(0)}{n^{m+1}}\,.$ (20)
+Let us start with the examination of (18) via “pointwise” approximations. We will use the same inequalities as in [6, Prop. 1]:
+$0\leq x$ $\implies 0\leq\mathsf{e}^{-m(1-x)}-x^{m}\,$,
+$\frac{1}{2}\leq x\leq 1$ $\implies 0\leq\mathsf{e}^{-m(1-x)}-x^{m}\leq m(1-x)^{2}\mathsf{e}^{-m(1-x)}\,.$ (21)
+Inequality (21), with $x=1-n^{-1}k$, $1\leq k\leq\frac{n}{2}$, is found in a similar context also in [14, answer to 6.3-34]. It suggests approximating $p_{m}(n)$ by
+$n^{-1}\sum_{0\leq k<n}\mathsf{e}^{-m(1-n^{-1}k)}=\frac{1-\mathsf{e}^{-m}}{n(\mathsf{e}^{n^{-1}m}-1)}\,$,
+which we rewrite as $(1-\mathsf{e}^{-m})q_{m}(n)$, defining for this purpose:
+$q_{m}(n)=\frac{1}{n(\mathsf{e}^{n^{-1}m}-1)}\,.$ (22)
+But the approximation of $p_{m}(n)$ by $q_{m}(n)$ becomes poor for $n$ large, as:
+$\lim_{n\to\infty}p_{m}(n)=\int_{0}^{1}x^{m}\mathrm{d}x=(m+1)^{-1},\text{ while }\lim_{n\to\infty}q_{m}(n)=m^{-1}.$
+Thus, we should either switch to $q_{m+1}(n)$ (we shall see later that it arises very naturally from Equation (20)), or approximate $(m+1)p_{m}(n)$ via $mq_{m}(n)$.
+We start with a Lemma which uses Equation (21) to compare $p_{m}(n)$ with $(1-\mathsf{e}^{-m})q_{m}(n)$ uniformly for the entire $n$ range.
+Lemma 1.
+There is a positive constant $K_{1}<3.1$ such that for every real positive $m$ and every positive integer $n$, there holds:
+$0\leq(1-\mathsf{e}^{-m})q_{m}(n)-p_{m}(n)\leq\frac{K_{1}}{m^{2}}\,.$ (23)
+Proof.
+We write:
+$(1-\mathsf{e}^{-m})q_{m}(n)-p_{m}(n)\begin{aligned} &=\Bigl(n^{-1}\sum_{0\leq k<n/2}+n^{-1}\sum_{n/2\leq k<n}\Bigr)\bigl(\mathsf{e}^{-m(1-n^{-1}k)}-(n^{-1}k)^{m}\bigr)\\ &=S_{1}+S_{2}\,.\end{aligned}$
+We bound $S_{1}$ (trivially) for $n$ even by $\int_{0}^{\frac{1}{2}}\mathsf{e}^{-m(1-x)}\mathrm{d}x=m^{-1}(\mathsf{e}^{-m/2}-\mathsf{e}^{-m})$ and for $n$ odd by $\int_{0}^{\frac{1}{2}-(2n)^{-1}}\mathsf{e}^{-m(1-x)}\mathrm{d}x+2\int_{\frac{1}{2}-(2n)^{-1}}^{\frac{1}{2}}\mathsf{e}^{-m(1-x)}\mathrm{d}x=m^{-1}\bigl(\mathsf{e}^{-m/2}-\mathsf{e}^{-m}+\mathsf{e}^{-m/2}(1-\mathsf{e}^{-m/(2n)})\bigr)$.
+This is maximal for $n=1$ so we adopt $2m^{-1}(\mathsf{e}^{-m/2}-\mathsf{e}^{-m})$ as general upper bound valid for all positive integers $n$ and positive real numbers $m$.
+Regarding $S_{2}$, we use the inequality (21), and after the change $k\to n-k$, and then extending the range of $k$ to infinity and letting $\tau=n^{-1}m>0$, we get
+$S_{2}\leq\sum_{1\leq k\leq n/2}mn^{-3}k^{2}\mathsf{e}^{-n^{-1}mk}\leq m^{-2}\sum_{k=1}^{\infty}\tau^{3}k^{2}\mathsf{e}^{-\tau k}=m^{-2}\tau^{3}\frac{\mathsf{e}^{-\tau}+\mathsf{e}^{-2\tau}}{(1-\mathsf{e}^{-\tau})^{3}}\,.$
+It turns out that this function of $\tau>0$ is nicely decreasing with its supremum attained for $\tau\to 0^{+}$ (this can be proven using the Poisson summation formula).
+So $m^{2}S_{2}<2$.
+Adding the bound for $S_{1}$ we obtain (23) with
+$K=2+\sup_{x>0}2x(\mathsf{e}^{-x/2}-\mathsf{e}^{-x}).$ (24)
+The supremum is attained at $x_{0}\approx\numprint{2.89115}$ (and values decrease for $x>x_{0}$) and is $\approx\numprint{1.04138}$.
+So we can take $K=\numprint{3.05}$.
+Restricting to $m\geq m_{0}\geq 3$, we get (23) with $K=2+2m_{0}(\mathsf{e}^{-m_{0}/2}-\mathsf{e}^{-m_{0}})$, and only $2+m_{0}(\mathsf{e}^{-m_{0}/2}-\mathsf{e}^{-m_{0}})$ for even $n$’s.
+∎
+As a corollary, if we let both $m$ and $n$ go to infinity such that their ratio converges to some value $q$, we get (with $1$ on the right if $q=0$):
+$\begin{split}\lim_{m,n\to\infty,mn^{-1}\to q}\frac{B_{m+1}(n)-B_{m+1}(0)}{n^{m+1}}=\lim\,(m+1)p_{m}(n)\\ =\lim\,mq_{m}(n)=\lim\frac{n^{-1}m}{\mathsf{e}^{n^{-1}m}-1}=\frac{q}{\mathsf{e}^{q}-1}\,.\end{split}$
+This is indeed what one would expect from expressing the Bernoulli polynomials in terms of Bernoulli numbers and taking the limit term-wise, but Lemma 1 makes it easy (and gives an $O(m^{-1})$ error estimate).
+Reference [15], which uses the saddle-point method in the complex domain, provides full asymptotic for, among others, $B_{n}(nz+\frac{1}{2})$ as $n\to\infty$.
+Lemma 2.
+There is a constant $K_{2}$ such that for every positive $m$:
+$0\leq\sum_{l=0}^{\lfloor\log_{2}(m+1)\rfloor}\Bigl((1-\mathsf{e}^{-m})q_{m}(2^{l})-p_{m}(2^{l})\Bigr)\leq\frac{K_{2}}{(m+1)^{2}}\,.$
+Proof.
+Let $L_{m}=\lfloor\log_{2}(m+1)\rfloor$.
+The $l=0$ term on the left-hand side of Equation (23) is, for $m>0$, $(1-\mathsf{e}^{-m})q_{m}(1)=(1-\mathsf{e}^{-m})(e^{m}-1)^{-1}=\mathsf{e}^{-m}\leq 4\mathsf{e}^{-1}(m+1)^{-2}$.
+We assume now $m\geq 1$, so that $L_{m}\geq 1$, and we need only consider $1\leq l\leq L_{m}$.
+Using the same notation as in the proof of Lemma 1, and Equation (21), we have, for any even positive integer $n$:
+$0\leq(1-\mathsf{e}^{-m})q_{m}(n)-p_{m}(n)\leq m^{-1}\mathsf{e}^{-m/2}+m^{-2}\sum_{k=1}^{\infty}\tau^{3}k^{2}\mathsf{e}^{-\tau k}\,$,
+with $\tau=n^{-1}m$.
+Summing for $2\leq 2^{l}\leq m+1$, the first term in the right-hand side contributes $L_{m}m^{-1}\mathsf{e}^{-m/2}$ which is a fortiori $O((m+1)^{-2})$.
+As per the second summands, as $\tau\geq m/(m+1)\geq\frac{1}{2}$, there holds $\mathsf{e}^{-\tau}\leq\mathsf{e}^{-\frac{1}{2}}$ and $\sum_{k\geq 1}k^{2}\mathsf{e}^{-\tau k}\leq\mathsf{e}^{-\tau}\sum_{k\geq 1}k^{2}\mathsf{e}^{-(k-1)/2}$.
+Hence, the second summand is, for each of $\tau=\frac{m}{2}$, $\frac{m}{4}$, …, $2^{-L_{m}}\geq\frac{1}{2}$, bounded by $Cm^{-2}\tau^{3}\mathsf{e}^{-\tau}$, for some constant $C$.
+We use the weaker upper bound $m^{-2}O(\mathsf{e}^{-\tau/2})$ and obtain in total $m^{-2}O(\mathsf{e}^{-1/4}+\mathsf{e}^{-1/2}+\dots+\mathsf{e}^{-2^{L_{m}}/8})=O(m^{-2})$.
+∎
+Let us set from here on, as in the previous proof:
+$L_{m}=\lfloor\log_{2}(m+1)\rfloor\,.$
+We also need an upper bound for $m\sum_{l=0}^{L_{m}}q_{m}(2^{l})=\sum_{l=0}^{L_{m}}\frac{m}{2^{l}(\mathsf{e}^{2^{-l}m}-1)}$.
+Not aiming at optimal results, we use the trivial bound valid for positive $x$:
+$\frac{x}{\mathsf{e}^{x}-1}=\frac{x\mathsf{e}^{-x/2}}{2\sinh\frac{x}{2}}<\mathsf{e}^{-x/2}.$
+Hence, as in the proof of Lemma 2, we get, supposing $m\geq 1$, hence $2^{L_{m}}\leq 2m$:
+$m\sum_{l=0}^{L_{m}}q_{m}(2^{l})$ $\leq\mathsf{e}^{-\frac{m}{2}}+\mathsf{e}^{-\frac{m}{4}}+\dots+\mathsf{e}^{-2^{-L_{m}}\frac{m}{2}}$
+$\leq\mathsf{e}^{-2^{L_{m}}/4}+\mathsf{e}^{-2^{L_{m}}/8}+\dots+\mathsf{e}^{-1/4}=O(1).$
+For $0<m<1$, there is only one term $\frac{m}{\mathsf{e}^{m}-1}=O(1)$. Hence:
+Lemma 3.
+There is a constant $K_{3}$ such that for every positive $m$:
+$0\leq\sum_{l=0}^{L_{m}}q_{m}(2^{l})\leq\frac{K_{3}}{m}\,.$
+Lemma 4.
+There is a constant $K_{4}$ such that for every positive $m$:
+$0\leq\sum_{l=0}^{\lfloor\log_{2}(m+1)\rfloor}\Bigl(q_{m}(2^{l})-p_{m}(2^{l})\Bigr)\leq\frac{K_{4}}{m^{2}}\,.$
+Proof.
+From Lemmas 2 and 3 the sum is bounded above by $K_{2}(m+1)^{-2}+\mathsf{e}^{-m}K_{3}m^{-1}$, so, multiplying by $m^{2}$ it is bounded by $K_{2}+\mathsf{e}^{-1}K_{3}$. ∎
+We now obtain:
+Lemma 5.
+There is a constant $K_{5}$ such that for every positive $m$:
+$\sum_{l=0}^{L_{m}}\Bigr|mq_{m}(2^{l})-(m+1)p_{m}(2^{l})\Bigl|\leq\frac{K_{5}}{m+1}$
+Proof.
+For $0<m<1$, we have only one term $mq_{m}(1)-(m+1)p_{m}(1)$ with $mq_{m}(1)=m(e^{m}-1)^{-1}\leq 1$, and $p_{m}(1)=0$, so it is bounded by 11 hence by $K(m+1)^{-1}$ for any $K\geq 2$.
+For $m\geq 1$, we first decompose
+$mq_{m}(2^{l})-(m+1)p_{m}(2^{l})=\begin{aligned} &(m+1)\Bigl((1-e^{-m})q_{m}(2^{l})-p_{m}(2^{l})\Bigr)\\ &-\bigl(1-(m+1)\mathsf{e}^{-m}\bigr)q_{m}(2^{l})\end{aligned},$
+hence,
+$\Bigl|mq_{m}(2^{l})-(m+1)p_{m}(2^{l})\Bigr|\leq(m+1)\Bigl((1-e^{-m})q_{m}(2^{l})-p_{m}(2^{l})\Bigr)+q_{m}(2^{l})\,.$
+Using Lemmas 2 and 3 we get, recalling $m\geq 1$,
+$\sum_{l=0}^{L_{m}}\Bigr|mq_{m}(2^{l})-(m+1)p_{m}(2^{l})\Bigl|\leq\frac{K_{2}}{m+1}+\frac{K_{3}}{m}\leq\frac{K_{2}+2K_{3}}{m+1}\,$,
+and the statement to be proven holds with $K_{5}=\max(2,K_{2}+2K_{3})$. ∎
+We now need to control the $2^{l}>m+1$ range. This is done via a completely different approach to the quantities $p_{m}(n)$ and $q_{m}(n)$ (or $q_{m+1}(n)$).
+Lemma 6.
+There are constants $K_{6}$ and $K_{7}$ such that
+$\sum_{l>L_{m}}\Bigr|mq_{m}(2^{l})-(m+1)p_{m}(2^{l})\Bigl|$ $\leq\frac{K_{6}}{m+1}$ (25)
+$\sum_{l>L_{m}}\Bigr|(m+1)q_{m+1}(2^{l})-(m+1)p_{m}(2^{l})\Bigl|$ $\leq\frac{K_{7}}{m+1}$ (26)
 Remark 6.
-A variant of (13) is (see Equation (22) for the Euler polynomials):
-$e_{m}$ $=\sum_{l=0}^{\infty}(l+1)\frac{B_{m+1}(2^{l+1})-2^{m+1}B_{m+1}(2^{l})+(2^{m+1}-1)B_{m+1}(0)}{2^{(l+1)(m+1)}}$
-$=\frac{m+1}{2}\sum_{l=0}^{\infty}(l+1)\frac{E_{m}(2^{l+1})-E_{m}(0)}{2^{(l+1)(m+1)}}\,.$
-One only needs to write the numerator as $(l+1)(v_{l+1}-2^{m+1}v_{l})$ with $v_{l}=B_{m+1}(2^{l})-B_{m+1}(0)-2^{(m+1)l}$ and rearrange the sum. Details are left to the reader.
-We will derive in this section from Equation (12) some asymptotic for $e_{m}$ as $m\to\infty$.
-From (11), the $l$th term is non-negative and bounded above by $(m+1)2^{-l-1}$.
-Let $L_{m}=\lfloor 2\log_{2}(m+1)\rfloor$.
-The sum of the contributions for $l>L_{m}$ is bounded above by $(m+1)2^{-L_{m}-1}$, hence, by $(m+1)^{-1}$.
-To approximate the ($l$ dependent) moments $p_{l}=2^{-l}\sum_{0\leq n<2^{l}}(2^{-l}n)^{m}$ for $l\leq L_{m}$, we shall apply the same technique as in [6, Prop.
-1].
-For the convenience of the reader, the presentation here is entirely self-contained.
-[^1] In v1 of [6], which is the current version at time of writing, the Proposition 1 is formulated in a way which does not make explicit the implied constants in the big-O, although this is quite easy to extract from the proof there.
-As here we need to apply this to mutliple probability measures, we had anyhow to reproduce parts of the argument from [6] in order to control the uniformity.
-Certainly, the published version of [6] will make more explicit the big-O bound.
-Let us define
-$q_{l}=2^{-l}\sum_{0\leq n<2^{l}}\mathsf{e}^{-m(1-2^{-l}n)}=\frac{1-\mathsf{e}^{-m}}{2^{l}(\mathsf{e}^{m2^{-l}}-1)}\,.$ (14)
-There holds $0\leq x\mathsf{e}^{-x}\leq\mathsf{e}^{-1}$ for $x\geq 0$, hence $x^{m}\leq\mathsf{e}^{-m(1-x)}$.
-In particular $(2^{-l}n)^{m}\leq\mathsf{e}^{-m(1-2^{-l}n)}$ and $0\leq p_{l}\leq q_{l}$.
-We split $q_{l}-p_{l}$ into two parts:
-$q_{l}-p_{l}=2^{-l}(\sum_{0\leq n<2^{l-1}}+\sum_{2^{l-1}\leq n<2^{l}})\biggl(\mathsf{e}^{-m(1-2^{-l}n)}-(2^{-l}n)^{m}\biggr)$ (15)
-We bound trivially the first part by
-$2^{-l}\sum_{0\leq n<2^{l-1}}\mathsf{e}^{-m(1-2^{-l}n)}\leq\int_{0}^{\frac{1}{2}}\mathsf{e}^{m(x-1)}\mathrm{d}x=\frac{1}{m}\mathsf{e}^{-m/2}(1-\mathsf{e}^{-m/2}).$
-This, and everything following applies also to non-integer positive $m$’s, the quantity $e_{m}$ can be defined for them using (12).
-Regarding the second part in Equation (15), we use (as in the proof of [6, Prop. 1]) the bound
-$m\geq 0,\;\frac{1}{2}\leq x\leq 1\implies\mathsf{e}^{-m(1-x)}-x^{m}\leq m(1-x)^{2}\mathsf{e}^{-m(1-x)}.$
-So, for positive $m$:
-$0\leq q_{l}-p_{l}\leq\frac{\mathsf{e}^{-m/2}}{m}+2^{-l}\sum_{2^{l-1}\leq n<2^{l}}m(1-2^{-l}n)^{2}\mathsf{e}^{-m(1-2^{-l}n)}$
-(Note that if $l=0$, the sum above is empty and has value 0; so we now assume $l\geq 1$ and the bound we obtain also applies with $l=0$).
-Let $\psi_{m}(x)=\frac{\mathrm{d}}{\mathrm{d}x}\left((1-x)^{2}\mathsf{e}^{-m(1-x)}\right)=\bigl(m(1-x)-2\bigr)(1-x)\mathsf{e}^{-m(1-x)}$.
-We write:
-$2^{-l}\sum_{2^{l-1}\leq n<2^{l}}m(1-2^{-l}n)^{2}\mathsf{e}^{-m(1-2^{-l}n)}=\frac{m}{8}\mathsf{e}^{-m/2}+m2^{-l}\sum_{2^{l-1}\leq n<2^{l}}\int_{\frac{1}{2}}^{2^{-l}n}\psi_{m}(y)\mathrm{d}y$
-Using the probability measure $\mu_{l}=2^{-l}\sum_{0\leq n<2^{l}}\delta_{2^{-l}n}$, which has the property $\mu_{l}([y,1])\leq 1-y$ for every $y\in[0,1]$, we can rewrite the second summand on the right-hand side as:
-$)\displaystyle m\int_{[\frac{1}{2},1\rparen}\int_{\frac{1}{2}}^{x}\psi_{m}(y)\mathrm{d}y\,\mu_{l}(\mathrm{d}x$) $\leq m\int_{\frac{1}{2}\leq y\leq x<1}\max(\psi_{m}(y),0)\mathrm{d}y\,\mu_{l}(\mathrm{d}x)$
-$\leq m\int_{\frac{1}{2}}^{1}\max(\psi_{m}(y),0)(1-y)\mathrm{d}y$
-$=m\int_{\frac{1}{2}}^{\max(1-\frac{2}{m},\frac{1}{2})}\bigl(m(1-y)-2\bigr)(1-y)^{2}\mathsf{e}^{-m(1-y)}\mathrm{d}y$
-$=m^{-2}\int_{\min(2,m/2)}^{m/2}(t-2)t^{2}\mathsf{e}^{-t}\mathrm{d}t\leq 18\mathsf{e}^{-2}m^{-2}.$
-We have obtained, for positive $m$ (with some loss of precision near zero, but we are only interested into $m\to\infty$):
-$0\leq q_{l}-p_{l}\leq(m^{-1}+\frac{m}{8})\mathsf{e}^{-m/2}+\frac{18}{\mathsf{e}^{2}m^{2}}\leq C_{1}m^{-2},$ (16)
-with $C_{1}=\max_{t\geq 0}(t+t^{3}/8)\mathsf{e}^{-t/2}+18\mathsf{e}^{-2}$.
-It turns out that the maximum is attained at $t\approx\numprint{5.04}$ and its value is $\approx\numprint{1.693}$ and $18\mathsf{e}^{-2}\approx\numprint{2.436}$ and $C_{1}<\numprint{4.13}$.
-Of course when $m$ is large $18\mathsf{e}^{-2}+\epsilon$ gives a tighter bound.
-The proof applies to any positive $m$, and for $m$ a positive integer it becomes a statement involving the values of Bernoulli polynomials at $N=2^{l}$.
-That $N$ was a power of $2$ is not important.
-We formulate here a dedicated statement, as we could not find such property in the standard references.
+In a certain way, Equation (26) is the one arising more naturally from the proof.
+But as we have already written up a proof for Lemma 5, which is expressed with the $q_{m}(2^{l})$’s, it is convenient to also prove Equation (25).
+Proof.
+Let $n$ be some positive integer. From Equation (20):
+$(m+1)p_{m}(n)=\frac{B_{m+1}(n)-B_{m+1}(0)}{n^{m+1}}=\sum_{k=0}^{m}\binom{m+1}{k}B_{k}n^{-k}\,.$
+We now approximate the binomial coefficients:
+$\binom{m+1}{k}$ $=\frac{1}{k!}(m+1)^{k}\prod_{0\leq p<k}(1-(m+1)^{-1}p)$
+$\implies 0$ $\leq\frac{(m+1)^{k}}{k!}-\binom{m+1}{k}\leq\frac{k(k-1)}{2(m+1)}\frac{(m+1)^{k}}{k!}\,.$
+We used $\prod_{0\leq p<k}(1-\epsilon_{p})\geq 1-\sum_{0\leq p\leq k}\epsilon_{p}$ which is valid for $0\leq\epsilon_{p}\leq 1$, $0\leq p<k$. Hence
+$\Bigl|\sum_{k=0}^{m}\frac{B_{k}}{k!}(m+1)^{k}n^{-k}-(m+1)p_{m}(n)\Bigr|\leq(m+1)^{-1}\sum_{k=0}^{m}\frac{|B_{k}|}{k!}\frac{k(k-1)}{2}\bigl((m+1)n^{-1}\bigr)^{k}$
+We assume from here on $n\geq m+1$, which is the case for the powers of $2$ we will consider, $n=2^{l}$, $l>L_{m}$.
+The finite sums in the previous inequality are then partial sums of convergent series.
+We replace k(k1)k(k-1) by the larger $k(k+1)$ which verifies $m+1\leq\frac{k(k+1)}{2}$ for $k\geq m+1$, $m>0$.
+This allows to extend our termwise estimates to the $k>m$ range.
+Thus, there holds:
+$\Bigl|\sum_{k=0}^{\infty}\frac{B_{k}}{k!}(m+1)^{k}n^{-k}-(m+1)p_{m}(n)\Bigr|\leq(m+1)^{-1}\sum_{k=2}^{\infty}\frac{|B_{k}|}{k!}\frac{k(k+1)}{2}\bigl((m+1)n^{-1}\bigr)^{k}.$
+We recognize in the left-hand side
+$\sum_{k=0}^{\infty}\frac{B_{k}}{k!}(m+1)^{k}n^{-k}=\frac{n^{-1}(m+1)}{\mathsf{e}^{n^{-1}(m+1)}-1}=(m+1)q_{m+1}(n)\,.$
+The series $S(\tau)=\sum_{k\geq 2}^{\infty}\frac{|B_{k}|}{k!}\frac{k(k+1)}{2}\tau^{k}$ has radius of convergence $2\pi$, so there is some constant $C$ such that $S(\tau)\leq C_{1}\tau^{2}$ for $0\leq\tau\leq 1$.
+Hence:
+$n\geq m+1\implies\Bigl|(m+1)q_{m+1}(n)-(m+1)p_{m}(n)\Bigr|\leq C_{1}(m+1)^{-1}\bigl((m+1)n^{-1}\bigr)^{2}.$
+For $n=2^{l}$, $l>L_{m}$, we have $n>(m+1)$ indeed. And summing $\bigl((m+1)n^{-1}\bigr)^{2}$ over such powers of $2$ is bounded above by $1+4^{-1}+16^{-1}+\dots$. Hence Equation (26).
+A modification in the handling the binomial coefficients will readily give Equation (25). We write this time:
+$\binom{m+1}{k}=\frac{1}{k!}m^{k}\prod_{-1\leq p<k-1}(1-m^{-1}p)\,.$
+The product (depending on $k\in[0,m]\cap\mathbb{Z}$) is bounded above by $1+m^{-1}$ and below by $\prod_{0\leq p<k-1}(1-m^{-1}p)$, hence the distance to $1$ is bounded in absolute value by
+$m^{-1}\delta_{k},\qquad\text{with }\delta_{k}=\max(1,\frac{(k-1)(k-2)}{2})\qquad(k\geq 1).$
+Moreover, $\delta_{0}=0$.
+We can use a weaker upper bound $k(k+1)/2$, so that $k(k+1)/2\geq m$ holds for $k\geq m+1$.
+This avoids having to discuss separately the (exponentially small, as one can check) contribution from the $k>m$ range.
+The sole differences with our earlier situation is that $k$ now starts at $1$ not at $2$ and that powers of $m+1$ are replaced by powers of $m$.
+And we obtain Equation (25).
+∎
 Proposition 3.
-For any two positive integers $m$ and $n$, the following uniform estimate holds:
-$0\leq\frac{1-\mathsf{e}^{-m}}{n\bigl(\mathsf{e}^{mn^{-1}}-1\bigr)}-\frac{B_{m+1}(n)-B_{m+1}(0)}{(m+1)n^{m+1}}\leq\frac{5}{m^{2}}.$ (17)
+Let, for positive real $m$,
+$f_{m}=\sum_{l=0}^{\infty}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)\,.$ (27)
+There holds $|e_{m}-f_{m}|\leq\frac{K}{m+1}$ for all $m>0$ and some constant $K$.
 Proof.
-Let’s use $N$ rather than $n$ which has served earlier as summation index.
-Recall $\sum_{0\leq n<N}n^{m}=(m+1)^{-1}\bigl(B_{m+1}(N)-B_{m+1}(0)\bigr)$.
-The proof leading to (16) works identically with an arbitrary positive integer $N$ in the place of $2^{l}$, if $N$ is even.
-If $N\geq 3$ is odd, $N=2M+1$, $M\geq 1$, we have to be careful about how we split (the analog of) $q_{l}-p_{l}$ into two parts.
-Formerly we compared the first part with an integral from 0 to $\frac{1}{2}$ but now, using $0\leq n<\frac{1}{2}N$, i.e., $0\leq n\leq M$, we see that the rectangle basis extends a bit further to $\frac{1}{2}+\frac{1}{2N}$.
-We could improve that (for $m$ large) by bounding the contribution of $x=M/N$ pointwise by $\mathsf{e}^{-m/2}$ (as done in [6, Prop.
-1] which handles general positive measures).
-But let us continue with $\int_{0}^{\frac{1}{2}+\frac{1}{2N}}\mathsf{e}^{m(x-1)}\mathrm{d}x<\mathsf{e}^{m/2N}m^{-1}\mathsf{e}^{-m/2}\leq\mathsf{e}^{m/6}m^{-1}\mathsf{e}^{-m/2}$, for the first part.
-For the second part, i.e.
-for the contributions of the points $x=N^{-1}n$, $M+1\leq n<N$, we can still use $x=\frac{1}{2}$ as reference point, there is no change, except that we could gain a little in the $\frac{m}{8}\mathsf{e}^{-m/2}$ term (as the weight of the half-interval is now not $1/2$, but $M/N$).
-Anyway, we end up with a slightly deteroriated final global upper bound $\mathsf{e}^{m/6}m^{-1}\mathsf{e}^{-m/2}+\frac{m}{8}\mathsf{e}^{-m/2}+18\mathsf{e}^{-2}m^{-2}$, which is $\leq C_{2}m^{-2}$ with $C_{2}=\max_{t\geq 0}(t\mathsf{e}^{-t/3}+\frac{t^{3}}{8}\mathsf{e}^{-t/2})+18\mathsf{e}^{-2}$.
-The maximum is again attained near $t=5$, it is found to be $\numprint{2.27}$ and $C_{2}<4.67$.
+The series defining $f_{m}$ has positive terms and is convergent.
+Lemmas 5 and 6 allow to bound $|e_{m}-f_{m}|=\Bigl|\sum_{l=0}^{\infty}\bigl(mq_{m}(2^{l})-(m+1)p_{m}(2^{l})\bigr)\Bigr|$, even after having moved the absolute values inside the summation.
 ∎
-As a corollary if we let both $m$ and $n$ go to infinity such that the ratio converges to some value $q$, we get (with $1$ on the right if $q=0$):
-$\lim_{m,n\to\infty,mn^{-1}\to q}\frac{B_{m+1}(n)-B_{m+1}(0)}{n^{m+1}}=\frac{q}{\mathsf{e}^{q}-1}\,.$
-This is indeed what one would expect from expressing the Bernoulli polynomials in terms of Bernoulli numbers and taking the limit term-wise, but (17) makes it easy (and gives an $O(m^{-1})$ error estimate).
-Reference [11], which uses the saddle-point method in the complex domain, provides full asymptotics for, among others, $B_{n}(nz+\frac{1}{2})$ as $n\to\infty$.
-We now resume the study of the quantities $e_{m}$. Recall that we have set $L_{m}=\lfloor 2\log_{2}(m+1)\rfloor$ and already estimated
-$0\leq\sum_{l>L_{m}}\bigl(1-(m+1)p_{l}\bigr)\leq\frac{1}{m+1},$
-so, combining this with Equations (12) and (16):
-$\sum_{0\leq l\leq L_{m}}\bigl(1-(m+1)q_{l}\bigr)\leq e_{m}\leq\sum_{0\leq l\leq L_{m}}\bigl(1-(m+1)q_{l}\bigr)+(L_{m}+1)(m+1)\frac{C_{1}}{m^{2}}+\frac{1}{m+1}\,.$
-Let us recall Equation (14) defining $q_{l}$ and record our current result:
-$e_{m}=\sum_{0\leq l\leq L_{m}}\bigl(1-(m+1)\frac{1-\mathsf{e}^{-m}}{2^{l}(\mathsf{e}^{m2^{-l}}-1)}\bigr)+O(\frac{\log m}{m}).$ (18)
-This motivates the definition of auxiliary coefficients $f_{m}$ and $g_{m}$, for positive real $m$:
-$f_{m}$ $=\sum_{l=0}^{\infty}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)$ (19)
-$g_{m}$ $=-\sum_{l=-1}^{-\infty}\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}$ (20)
-Clearly the series defining $f_{m}$ converges uniformly for $m$ bounded. We bound trivially the terms in the series defining $g_{m}$::
+6. Interlude on another sequence
+As part of [14, Ex. 6.3-34] the following is asked (we have translated to our notation). Let for $m\geq 2$
+$a_{m}$ $=\sum_{l=1}^{\infty}\biggl(\sum_{1\leq k<2^{l}}(2^{-l}k)^{m-1}-\frac{2^{l}}{m}+\frac{1}{2}\biggr)$
+$b_{m}$ $=\sum_{l=1}^{\infty}\biggl(\frac{1}{\mathsf{e}^{2^{-l}m}-1}-\frac{2^{l}}{m}+\frac{1}{2}\biggr).$
+Show $a_{m}=b_{m}+O_{m\to\infty}(m^{-1})$.
+We note that the general term $a_{m}(l)$ in the $a_{m}$ series is $O(2^{-l})$.
+This follows from Euler-Maclaurin, or directly from using Bernoulli polynomials:
+$a_{m}(l)=\frac{B_{m}(2^{l})-B_{m}(0)}{m2^{l(m-1)}}-\frac{2^{l}}{m}+\frac{1}{2}=\sum_{k=2}^{m-1}\binom{m}{k}m^{-k}B_{k}m^{k-1}(2^{-l})^{k-1}\,.$ (28)
+We used $B_{0}=1$, $B_{1}=-\frac{1}{2}$ and isolated a term $m^{-k}$ to merge it with the binomial coefficient. Similarly to the proof of Lemma 6, we now take note of the inequality
+$0\leq\frac{1}{k!}-\frac{\binom{m}{k}}{m^{k}}\leq\frac{1}{k!}\frac{k(k-1)}{2m}\,.$
+Let $b_{m}(l)$ be the $l$th term in $b_{m}$.
+We see that we are going to obtain $\sum_{2^{l}>m}|b_{m}(l)-a_{m}(l)|=O(m^{-1})$ via the similar steps as in the proof of Lemma 6.
+It is important that in Equation (28), the smallest exponent of $m2^{-l}$ is $1>0$.
+There remains to evaluate the contributions for $1\leq 2^{l}\leq m$. Using the notation from the previous section, it is a question of upper bounding
+$\sum_{1\leq 2^{l}\leq m}\left|2^{l}p_{m-1}(2^{l})-2^{l}q_{m}(2^{l})\right|.$
+This is at most:
+$S_{1}+S_{2}=\sum_{1\leq 2^{l}\leq m}\left|2^{l}p_{m-1}(2^{l})-2^{l}q_{m-1}(2^{l})\right|+\sum_{1\leq 2^{l}\leq m}\left|2^{l}q_{m-1}(2^{l})-2^{l}q_{m}(2^{l})\right|\,.$
+For $S_{1}$, we use $2^{l}\leq m$ and Lemma 4 to see that $S_{1}\leq mK_{4}(m-1)^{-2}=O(1/m)$. Regarding $S_{2}$, we estimate, for each positive integer $n$:
+$(\mathsf{e}^{n^{-1}(m-1)}-1)^{-1}-(\mathsf{e}^{n^{-1}m}-1)^{-1}$ $=\frac{\mathsf{e}^{n^{-1}\xi}}{n(\mathsf{e}^{n^{-1}\xi}-1)^{2}}\quad(\xi\in[m-1,m])$
+$\leq\frac{1}{n4\sinh^{2}((2n)^{-1}\xi)}\leq\frac{n}{\xi^{2}}\leq\frac{n}{(m-1)^{2}}\,.$
+So $S_{2}\leq 2m/(m-1)^{2}=O(1/m)$.
+Remark 7.
+As we mentioned already, Theorem 4.1 from [17] (which probably implies the above $O(1/m)$ in view of its many parameters) has an invalid proof.
+It is apparent that our way of handling binomial coefficients should fix this issue, in general.
+7. Asymptotics
+Clearly the series from Equation (27) defining $f_{m}$ converges uniformly for $m$ bounded (for now, $m$ is real positive). We define (at first, for positive real $m$):
+$g_{m}=-\sum_{l=-1}^{-\infty}\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\,.$ (29)
+We bound trivially the terms in the series defining $g_{m}$:
 $x>0\implies\frac{x}{\mathsf{e}^{x}-1}=\frac{x\mathsf{e}^{-x/2}}{2\sinh(x/2)}\leq\mathsf{e}^{-x/2}$
-$0\leq-g_{m}\leq\frac{2m}{\mathsf{e}^{2m}-1}+\sum_{n=4}^{\infty}\mathsf{e}^{-nm/2}=O_{m\to\infty}(me^{-2m}).$
+$0\leq-g_{m}\leq\frac{2m}{\mathsf{e}^{2m}-1}+\sum_{n=4}^{\infty}\mathsf{e}^{-nm/2}=O_{m\to\infty}(me^{-2m}).$ (30)
 […]
@@
 Details are left to reader: $\sum_{l=0}^{\infty}(1-\frac{2^{-l}z}{\mathsf{e}^{2^{-l}z}-1})$ defines a meromorphic function in the entire complex plane, with poles at the elements of $2\pi i\mathbb{Z}\setminus\{0\}$, so in particular if we substitute $z=2^{t}$, we obtain an analytic function in the strip $|\Im t|<\frac{\pi}{2\log 2}$.
-Regarding $\sum_{k=1}^{\infty}\frac{2^{k}2^{t}}{\mathsf{e}^{2^{k}2^{t}}-1}$, it is absolutely and uniformly convergent if $a<\Re t<b$, $|\Im t|<\frac{\pi}{2}-\epsilon$, $0<\epsilon<\frac{\pi}{2}$, from bounding below the denominator and above the numerator.
+Regarding $\sum_{k=1}^{\infty}\frac{2^{k}2^{t}}{\mathsf{e}^{2^{k}2^{t}}-1}$, it is absolutely and uniformly convergent if $a<\Re t<b$, $|\Im t|<\frac{\pi}{2\log(2)}(1-\epsilon)$, $0<\epsilon<1$ (so that $\Re 2^{t}>2^{a}\sin\frac{\pi\epsilon}{2}$), from bounding below the denominator and above the numerator.
 This proves that $\phi(t)$ is as stated.
 ∎
-Returning to Equation (18), we get:
-$e_{m}$ $=L_{m}+1-\frac{m+1}{m}(1-\mathsf{e}^{-m})\biggl(L_{m}+1-\sum_{0\leq l\leq L_{m}}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)\biggr)+O(\frac{\log m}{m})$
-$=(1+O(\frac{1}{m}))\Bigl(f_{m}-\sum_{l>L_{m}}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)\Bigr)+O(\frac{\log m}{m}).$
-For $l>L_{m}=\lfloor 2\log_{2}(m+1)\rfloor$, $2^{-l}m\leq m/(m+1)^{2}$, hence:
-$\sum_{l>L_{m}}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)=O(\sum_{l>L_{m}}\frac{m^{2}}{4^{l}})=O(\frac{m^{2}}{4^{L_{m}}})=O(m^{-2}).$
-This gives, for $m\to+\infty$:
-$e_{m}=(1+O(\frac{1}{m}))\Bigl(\log_{2}(m)+\frac{1}{2}+\phi(\log_{2}(m))+O(m^{-2})+O(m\mathsf{e}^{-2m})\Bigr)+O(\frac{\log m}{m}).$
-As $\phi$ is periodic, it is bounded and we have obtained the proof of the main estimate as given in Theorem 2:
+Since $\phi(\log_{2}(m))=-\log_{2}(m)-\frac{1}{2}+f_{m}+g_{m}$, we now get from Proposition 3 and the bound (30):
 Proposition 5.
 The quantities $(e_{m})$ defined by the recurrence (1), or, more generally, defined for all real positive $m$ by Equation (15), verify the asymptotic estimate, as $m\to+\infty$:
-$e_{m}=\log_{2}(m)+\frac{1}{2}+\phi(\log_{2}(m))+O(\frac{\log m}{m}),$
+$e_{m}=\log_{2}(m)+\frac{1}{2}+\phi(\log_{2}(m))+O(\frac{1}{m}),$
 where $\phi(t)$ is the function defined by Equation (31). It is 11-periodic and analytic for $|\Im t|<\frac{\pi}{2\log(2)}$.
@@
 We have used the basic Riemann formula $\Gamma(s)\zeta(s)=\int_{0}^{\infty}x^{s-1}(\mathsf{e}^{x}-1)^{-1}\mathrm{d}x$ for $\Re s>1$.
-Using the usual bound $|\zeta(1+it)|=O_{t\to+\infty}(\log t)$ ([12, Thm.
+Using the usual bound $|\zeta(1+it)|=O_{t\to\pm\infty}(\log|t|)$ ([18, Thm.
 3.5]), and $|\Gamma(1+it)|\sim\sqrt{2\pi|t|}\exp(-\frac{\pi}{2}|t|)$ from Stirling formula in the complex domain (or more expediently from the reflection formula $\Gamma(s)\Gamma(1-s)=\pi/\sin(\pi s)$), we recover the analyticity of $\phi(t)$ in the strip $|\Im t|<\pi/\log(4)$ which was stated in Proposition 4.
-This completes the proof of Theorem 2.
+This is one among the Fourier series obtained by Szpankowski [17, Eq. (4.20)] and by Prodinger [16, Thm. 7]. The proof of Theorem 2 is complete.
 9. Exponential generating function
@@
 $f_{m}=\sum_{l=1}^{\infty}\frac{l2^{-l}m}{\mathsf{e}^{2^{-l}m}+1}\,.$
-The author did not find the next formula in standard references [1, 7, 8] although it is undoubtly well-known, and is related to the relation
+For the proof we will use the Lemma 7 below, which is undoubtedly very well-know, but which we did not locate in standard references [1, 7, 8]. It also relates to the relation
 $E_{n-1}(x)=2n^{-1}\bigl(B_{n}(x)-2^{n}B_{n}(2^{-1}x)\bigr)$ (32)
@@
 $1+(\mathsf{e}^{z}-1)^{-1}=\sum_{k=1}^{\infty}\bigl(2^{-k}-\frac{2^{-k}}{\mathsf{e}^{z/2^{k}}+1}\bigr)+z^{-1},$
-which, after canceling out the $1$ and multiplying by $z$ gives Equation (23). ∎
+which, after canceling out the $1$ and multiplying by $z$ gives Equation (33).
+Another method starts from the identity $\sinh w=w\cosh(\frac{w}{2})\cosh(\frac{w}{4})\cosh(\frac{w}{8})\cdots$ and computes the logarithmic derivative at $w=\frac{z}{2}$.
+∎
 Proof of Proposition 6.
@@
 Remark 8.
-Hence, if we let $E(t)$ denote the left-hand side of Equation (24), we see that $E(m)=(\mathsf{e}^{m}-1)f_{m}=(\mathsf{e}^{m}-1)\bigl(e_{m}+O(\frac{\log m}{m})\bigr)=\mathsf{e}^{m}e_{m}\bigl(1+O(m^{-1})\bigr)$ and thus $e_{m}=(1+O(m^{-1}))\mathsf{e}^{-m}E(m)=\mathsf{e}^{-m}E(m)+O(\frac{\log m}{m})$.
-This is very reminiscent of the general estimate from [6, Prop.
+Hence, if we let $E(t)$ denote the left-hand side of Equation (34), $E(m)=(\mathsf{e}^{m}-1)f_{m}=\mathsf{e}^{m}f_{m}+O(\log m)$, $f_{m}=\mathsf{e}^{-m}E(m)+O(\log(m)\mathsf{e}^{-m})$, $e_{m}=\mathsf{e}^{-m}E_{m}+O(m^{-1})$.
+This is very reminiscent of the estimate from [6, Prop.
 1] relating moments of some measures on the unit interval with their exponential generating function.
+It seems that such relations are a well-known theme for the asymtotics of sequences arising from the theoretical analysis of search algorithms, see [14, 9].
 Remark 9.
@@
 Comparison with Equation (31) shows that the above is exactly $\phi(\operatorname{Log}_{2}z)$.
-And indeed, the image of the right half-plane under $z\mapsto\operatorname{Log}_{2}z$ is the horizontal strip where we have defined $\phi$ as an nalytic function.
+And indeed, the image of the right half-plane under $z\mapsto\operatorname{Log}_{2}z$ is the horizontal strip where we have defined $\phi$ as an analytic function.
 Whether this relation between $\Psi$ and $\phi$ is indicative of some closer connection between the sequence $(e_{m})$ and the Thue-Morse sequence [2, 3] is currently unknown to the author.
+Acknowledgements. Thanks to Jean-Paul Allouche for drawing the author attention to A372422 and to the work by Knuth and Prodinger.